### v1

# latex‘h’ float specifier changed to ‘ht’

Abstract.
We show that for the standard map family, for all values of the parameter, except one, the mapping has positive topological entropy. The main tool is the following result.
Let $S$ be a compact connected orientable surface and $f:S\rightarrow S$ an area preserving orientation preserving $C^{1}$ diffeomorphism of $S$. Assume that $U$ is an invariant domain of $S$ such that $fr_{S}{U}$ has a finite number of connected components.
Let $b$ be a regular ideal boundary point of $U$ which is fixed under the induced action by $f$ on the ideal boundary of $U$, and let $\hat{f}:C(b)\rightarrow C(b)$ the homeomorphism on the corresponding circle of prime ends.
Let $Z(b)$ be the impression of $b$ in $S$ and assume that all fixed points of $f$ in $Z(b)$ are non degenerate.
If there exists a fixed prime end $e\in C(b)$ then we know the following.
  1. (1)
    If $p$ is the principal point of $e$ then $p$ is also a fixed point of $Z(b)$ and $p$ is a saddle.
  2. (2)
    $C(b)$ has a finite number of fixed prime ends and there exists a finite singular covering $\phi:C(b)\rightarrow Z(b)$, which is a semiconjugacy between the mapping of prime ends on $C(b)$ and the restriction of $f$ to $Z(b)$. In particular, $Z(b)$ is the connected union of finitely many saddle connections and the corresponding saddles.
This can be seen as a two dimensional generalization of the dynamics of homeomorphisms of the circle with fixed points.
2020 Mathematics Subject Classification:
Primary: 37; Secondary: A10, B40, B20, C29, E30.
Universidade Federal de Minas Gerais, Belo Horizonte, Brasil.
fernando.oliveira.f@gmail.com
Keywords: Standard map, homoclinic, prime end, ideal boundary, boundary dynamics.
1. Introduction
1.1. Statement of results
The standard map is a one parameter family of area preserving diffeomorphisms of the two dimensional torus $T^{2}=\mathbb{R}^{2}/\mathbb{Z}^{2}$ given by
$f_{\mu}(x,y)=(x+y+\frac{\mu}{2\pi}\sin(2\pi x),y+\frac{\mu}{2\pi}\sin(2\pi x))$, $\mu\in\mathbb{R}$.
This map is a model for many physical phenomena. It describes Poincaré’s surface of section of the kicked rotator, which consists of a stick that is free from gravitational forces, that rotates without friction around an axis located at one of its ends, and is periodically kicked at the other end. The kicked rotator approximates systems studied in many fields of physics. For example, circular particle accelerators accelerate particles by applying periodic kicks, as they circulate in the beam tube. Therefore, the structure of the trajectory of the beam can be approximated by the kicked rotator. For a more detailed account of the development of the problem, Lazutkin’s work on it, and results on the existence of transverse homoclinic orbits for small values of the parameter, see [7] and [8].
$f_{0}$ is just a linear twist and $f_{\mu}$ and $f_{-\mu}$ are conjugated. So we only consider parameters $\mu>0$.
For $\mu\neq 0$ there are two fixed points, $p=(0,0)$ and $q=(\frac{1}{2},0)$. $p$ is always a saddle with positive eigenvalues and it is called the principal fixed point of $f_{\mu}$. $q$ is elliptic if $40<\mu<$ 4 and a hyperbolic fixed point with negative eigenvalues if $\mu>4$.
One of our main results is the following.
Theorem If $\mu\neq 4$ then the standard map satisfies the following:
  1. (1)
    The four branches of $p=(0,0)$ have topologically transverse homoclinic points.
  2. (2)
    $f_{\mu}$ has positive topological entropy.
  3. (3)
    There exist periodic saddles with transverse homoclinic points.
This is Theorem 5.1 and section $5$ is all dedicate to its proof, which is based on the ideas developed in sections $2$, $3$ and $4$, and the use of the Inverse Function Theorem applied to a neighborhood of $p$.
In section $2$ we develop the basic ideas of the theories of the ideal completion of a surface $S$ and the prime ends compactification of a domain $U\subset S$. The main idea is to add to $U$ a circle of prime ends corresponding to each ”nontrivial” ideal boundary point of $U$, as a surface contained in $S$.
Ideal boundary points of a surface $S$ are also known as the topological ends of $S$, and we denote them by $b(S)$. The structure of $b(S)$ and the ideal completion of $S$, $B(S)=S\sqcup b(S)$, are nicely described in Proposition 2.9. $B(S)$ is almost a surface, and except for some elements of $b(U)$ that are accumulated by handles or Mobius bands, $B(S)$ is locally a surface.
If $U$ is a domain of $S$, then every ideal boundary point $b\in b(U)$ leaves an impression in $S$, like a mark. The set
$Z(b):=\{x\in S\mid\exists(x_{n})\text{ in }U\text{ with }{x_{n}}\rightarrow b\text{ in }B(U)\text{ and }x_{n}\rightarrow x\text{ in }S\}$
is called the impression of $b$ in $S$. Roughly, it tells the accumulation in $S$ of sequences $(x_{n})$ in $U$ such that converge to $b$ in $B(U)$.
Obviously $Z(b)$ is a subset of the frontier of $U$ in $S$, $fr_{S}U$.
We will be interested in replacing impressions by circles of prime ends, in the case where the impression $Z(b)$ is a compact subset of $S$ with more than one point. Is this case $b$ is called a regular ideal boundary point, and denote these set of such points by $b_{reg}(U)$.
If $b\in b(U)$ and $Z(b)$ is a one point set contained in $S$, then $b$ corresponds to a puncture of $U$ in $S$, and we replace $Z(b)$ by one point.
We will work with domains $U$ such that $fr_{S}U$ is compact and has a finite number of connected components. In this case there exist only a finite number of ideal boundary points $b\in b(U)$ such that $Z(b)$ is compact. This follows from the following result (Theorem 2.15).
Theorem Let $K$ be a compact subset of a compact connected surface $S$ and $U$ a residual domain of $K$. If $K$ has $m$ connected components then $U$ has at most $m(g+1)$ ideal boundary points, where $g$ is the genus of $S$.
With this result applied to $K=fr_{S}U$, we can give an elementary proof of Lemma $2.3$ of [13], providing an upper bound for the number of regular ideal boundary points of $U$.
If $S$ is a connected surface, $U$ is a domain of $S$ such that $fr_{S}U$ is compact and has a finite number of connected components, then $b_{reg}(U)$ is a finite set. $U$ could have infinite genus and any orientability type. In this context, it is possible to defined the prime ends compactification $E(U)$ of $U$. For each $b\in b_{reg}(U)$ we replace $Z(b)$ by a circle of prime ends and obtain a surface with compact boundary $\hat{U}$ and $E(U)$ is homeomorphic to $B(\hat{U})$. The definition of $E(U)$ is the same as the one given by Mather in [14], except that we add the condition that the end cuts be pairwise disjoint compact subsets of $S$. All of Mather’s results extend this more general context. These results have already been written in the context homeomorphisms and domains of infinite genus, but this general approach made the article too long.
Therefore we chose to simplify and assume compactness and smoothness of maps in order to focus on some important ideas. More general results will appear later.
Let $S$ be a compact surface and $U$ a domain of $S$ such that $fr_{S}U$ has a finite number of components. It follows that $b(U)$ is a finite set (see Theorem 2.14).
In this situation it is possible to define the prime ends compactification of $U$ following [14].
We replace the impression $Z(b)$ of each $b\in b(U)$ with a circle of prime ends $C(b)$ or a point, depending on whether $b$ is regular or not. In this way, we obtain a compact surface with boundary $E(U)$ whose boundary components are the circles of prime ends $C(b)$, where $b$ ranges over all points of the set of regular ideal boundary points of $U$.
Each prime end represents a way of approaching points of $Z(b)\subset fr_{S}U$, as we move out of $U$ in the direction of $b$.
Firstly consider the simplest case where $U$ is homeomorphic to an open disk. In this case $b(U)$ consists of one ideal boundary point, $Z(b)=fr_{S}U$, $B(U)$ is homeomorphic to a sphere and $E(U)$ is homeomorphic to a closed disk.
We would like to start a description of the other main result of this paper, which is in section 3.
We study area preserving diffeomorphisms of compact surfaces $f:S\rightarrow S$ and the frontier $fr_{S}U$ of invariant domains $U$, such that $fr_{S}U$ has a finite number of components. The main problem is to describe the structure of $fr_{S}U$ and the dynamics of $f$ restricted to $fr_{S}U$.
We have the following result (Corollary 3.14).
Proposition Let $S$ be a compact connected orientable surface and $f:S\rightarrow S$ an area preserving orientation preserving $C^{1}$ diffeomorphism of $S$.
Let $U$ be an invariant set homeomorphic to an open disk, and suppose that all fixed points of $f$ in $fr_{S}U$ are non degenerate.
If there exists a fixed prime end $e\in C(b)$ then $fr_{S}U$ is a connected finite union of connections and the corresponding saddle fixed points.
Each invariant manifold of a saddle has two branches. A connection is a branch contained in two invariant manifolds. It connects two saddles (that may coincide).
Our result can be seen as a generalization to dimension $2$ of Poincaré’s result about the dynamics of orientation preserving circle homeomorphisms with fixed points. In this case all periodic points are fixed and the arcs between fixed points are connections in the sense that if $x$ belongs to one of these arcs, then $lim_{n\rightarrow\pm\infty}{f^{n}(x)}$ are the end points points of the arc.
Of course we have the same results if the homeomorphism on the circle of prime ends has rational rotation number.
Before we proceed, note that a homeomorphism $f:S\rightarrow S$ induces homeomorphisms $f_{*}:B(U)\rightarrow B(U)$ and $\hat{f}:E(U)\rightarrow E(U)$.
The second main result of this paper is the following (Theorem 3.1).
Theorem Let $S$ be a compact connected orientable surface and $f:S\rightarrow S$ an area preserving orientation preserving $C^{1}$ diffeomorphism of $S$. Assume that $U$ is an invariant domain of $S$ such that $fr_{S}{U}$ has a finite number of connected components.
Let $b$ be a regular ideal boundary point of $U$ such that $f_{*}(b)=b$ and $\hat{f}:C(b)\rightarrow C(b)$ the homeomorphism on the corresponding circle of prime ends. Suppose that all fixed points of $f$ in $Z(b)$ are non degenerate.
Assume that there exists a fixed prime end $e\in C(b)$. Then we know the following.
  1. (1)
    If $p$ is the principal point of $e$ then $p$ is also a fixed point of $Z(b)$ and $p$ is a saddle.
  2. (2)
    $C(b)$ has a finite number of fixed prime ends and there exists a finite singular covering $\phi:C(b)\rightarrow Z(b)$, which is a semiconjugacy between the mapping of prime ends on $C(b)$ and the restriction of $f$ to $Z(b)$. In particular $Z(b)$ is the connected union of finitely many saddle connections and the corresponding saddles.
This is a result about only one circle of prime ends. If all circles of prime ends of $U$ have a fixed point (or more generally, rational rotation number), then the whole frontier of $U$ is a finite union of connections.
Many articles, which deal with related results, use the hypothesis that elliptic periodic points that belong to $fr_{S}U$ and arise as principal points of periodic prime ends are Moser stable. This was the way to get rid of them, which is no longer necessary: the principal point of a fixed prime end can not be elliptic.
In section $4$ we study recurrence and the existence of homoclinic points for area preserving diffeomorphisms of compact surfaces. The main tool is Theorem 3.1.
We prove that if $L$ is an invariant branch of a saddle $p$ and all fixed points of $f$ contained in $cl_{S}L$ are non degenerate, then either $L$ is a connection or $L$ accumulates on both sectors adjacent to itself.
It follows that in the absence of connections the self accumulation of branches brings complicated dynamics, and it is possible to prove the following result about the existence of homoclinic points in lower genus (Propositions 4.2 and 4.3).
Proposition Let $S$ be the two dimensional sphere $S^{2}$ or the torus $T^{2}$ and $f:S\rightarrow S$ an area preserving orientation preserving $C^{1}$ diffeomorphism of $S$.
Let $p$ be a saddle fixed point of $f$ and suppose that the branches of $p$ are invariant sets that are not connections. Assume that all fixed points of $f$ contained in $cl_{S}(W_{p}^{u}\cup W_{p}^{s})$ are non degenerate.
Then all pairs of adjacent branches of $p$ intersect.
Note that this proposition and the entire article only deal with fixed points. We are not allowed to freely take powers of diffeomorphisms. We try to make all hypotheses as weak as possible. This makes results more easily applicable to specific examples, like the Standard Map.
This last proposition is used in the proof of the Theorem 5.1.
1.2. Notation
In this article, $S$ will denote a smooth connected boundaryless surface, provided with a measure $\mu$, which is finite on compact sets and positive on open sets. We say that a homeomorphism $f:S\rightarrow S$ is measure preserving if $\mu\big{(}f^{-1}(E)\big{)}=\mu(E)$ for every Borel subset $E$ of $S$.
From a certain point onwards we will assume hypotheses such as the compactness of $S$, its orientability, that $f$ is a $C^{1}$ diffeomorphism and that $f^{*}\eta=\eta$, for some non degenerate $2$-form on $S$. In this case, we will still denote by $\mu$ the invariant measure induced by $\eta$.
By a domain we mean an open connected subset of $S$. If $A\subset B$, we use the notation $int_{B}(A)$, $cl_{B}(A)$ and $fr_{B}(A)$ for the interior, closure and the frontier of $A$ in $B$, respectively. By a disk (open, closed) we mean a set homeomorphic to the disk (open, closed) $D$ in the Euclidean norm of $\mathbb{R}^{2}$, and by a circle the analogue. If $K$ is a compact subset of $S$, a connected component of $S\setminus K$ is called a residual domain of $K$.
Most of the time, we refer to the connected components of a topological space simply as components. When referring to an ideal boundary component, the three words will always appear together.
1.3. A note on earlier announcements of Theorem 3.1
Theorem 3.1 was announced by this author in 2009/2010 at several conferences and seminars in different countries, including the International Conference on Dynamical Systems – 2010, Celebrating the 70th birthday of Jacob Palis, Búzios, Rio de Janeiro, from 02/25 to 03/05, 2010. See [9] for information about speakers and titles of their talks at the Conference. Unfortunately, for family reasons, it is only now that we have been able to write the results.
In 2018, this result was announced again in [11]. In their announcement they mention my previous announcement as ”a similar result”. See Theorem 1.4 and the paragraph following it in [11].
They state Theorem 3.1 with slightly different assumptions but exactly the same thesis.
Their result is stated only for domains of $\mathbb{R}^{2}$ homeomorphic to an open disk, and the generalizations they claim to have are only for maps homotopic to the identity (see the last paragraph of Section 2.4 of [11]).
On the other hand, we have no restrictions on the action on homology.
They are never clear about what kind of surfaces and periodic points they can handle.
There is one important point: as far as we know, so far they have not published the result and have never shown a complete argument or even a preprint.
2. Topological preliminaries
2.1. The ideal completion of a surface
We are going to describe a compactification of $S$ by the addition of its ends or ideal boundary points. See [1], [12], [13], and [15].
We start by considering decreasing sequences $(P_{n})_{n\geq k}$ of non-empty open connected subsets of $S$ ($P_{n}\supset P_{n+1}\>\forall\>n\geq k)$ such that $fr_{S}P_{n}$ is compact for every $n\geq k$.
Let $(C_{n})$ be a decreasing sequence of subsets of $S$. We say that $(C_{n})$ leaves compact subsets of $S$ if for every compact subset $K$ of $S$ there exists $n_{0}$ such that $K\cap C_{n}=\emptyset$ for $n\geq n_{0}$.
Let $(P_{n})$ be a decreasing sequence of non-empty open connected subsets of a surface $S$ such that $fr_{S}P_{n}$ is compact for every $n$. It is easy to see that the following conditions are equivalent:
  • $(P_{n})$ leaves compact subsets of $S$
  • $\cap_{n}cl_{S}P_{n}=\emptyset$.
2.1 Definition.
An ideal boundary component of $S$ (IBC) is a decreasing sequence $(P_{n})$ of non-empty open connected subsets of $S$ such that $fr_{S}P_{n}$ is compact for every $n$ and $(P_{n})$ satisfies one of the two equivalent conditions above.
The following example shows that there exist decreasing sequences of open connected subsets $(P_{n})$ such that $fr_{S}P_{n}$ is compact for every $n$ and $\cap_{n}P_{n}=\emptyset$, but $\cap_{n}cl_{S}P_{n}\neq\emptyset$ and $(P_{n})$ does not leave compact subsets of $S$. This shows that we can not weaken the condition $\cap_{n}cl_{S}P_{n}=\emptyset$ in the definition of IBCs to $\cap_{n}P_{n}=\emptyset$.
2.2 Example.
[…]
Let $(P_{n})$ and $(Q_{m})$ be ideal boundary components of $S$. We say that $(P_{n})$ and $(Q_{m})$ are equivalent if for every $n$ there is $m$ such that $P_{n}\subset Q_{m}$ and vice versa. An end or ideal boundary point of $S$ is an equivalence class of ideal boundary components. The set of ideal boundary points, denoted by $b(S)$, is called the ideal boundary or the set of ends of $S$, and the disjoint union $B(S):=S\sqcup b(S)$ is called the ideal completion or the ends compactification of $S$.
Let $V$ be an open subset of $S$. We define $V^{\prime}$ as the set of ideal boundary points of $S$ whose representing IBCs $(P_{n})$ satisfy $P_{n}\subset V$ for some $n$. Let $V^{*}:=V\sqcup V^{\prime}$. Obviously $V^{\prime}\cap W^{\prime}=(V\cap W)^{\prime}$ and $V^{*}\cap W^{*}=(V\cap W)^{*}$. On $b(S)$ and $B(S)$ we are going to consider the topologies generated by the bases of sets formed by $V^{\prime}$ and $V^{*}$, respectively, where $V$ runs through the collection of all connected open subsets of $S$ with $fr_{S}V$ compact.
We are going to write $b=[(P_{n})]$ or just $b=(P_{n})$ to say that $(P_{n})$ is an IBC representing $b\in b(S)$.
A characterization of $B(S)$ as a compactification of $S$ is given by the following result:
2.3 Proposition.
$B(S)$ is a compactification of $S$ that satisfies the following properties:
  1. (1)
    $B(S)$ is a locally connected Hausdorff space.
  2. (2)
    $b(S)$ is totally disconnected.
  3. (3)
    $b(S)$ is non separating on $B(S)$ (meaning that for any open connected subset $V$ of $B(S)$, $V\setminus b(S)$ is connected).
If $M$ is another compactification of $S$ that satisfies these properties, then there exists a homeomorphism from $M$ onto $B(S)$ which is the identity on $U$.
For a proof see sections 36 and 37 of chapter 1 of [1].
2.1.1. Exhaustions
Now we would like to describe a practical method for calculating ideal boundary points.
An exhaustion of $S$ is an increasing sequence $(F_{n})$ of compact connected bordered surfaces contained in $S$, such that for every $n$ we have the following:
  • $F_{n}\subset F_{n+1}^{\circ}=int_{S}{F_{n+1}}$.
  • $S=\cup_{n}F_{n}$.
  • If $W$ is a connected component of $S\setminus F_{n}$, then $cl_{S}W$ is a non compact bordered surface whose boundary consists of exactly one connected component of $\partial F_{n}$ (one circle).
We have the following existence result.
2.4 Proposition.
Every non compact connected surface admits an exhaustion.
For a proof see Theorem of section 29A of chapter 1 of [1]. Now we would like to state another well known result.
2.5 Proposition.
Let $(F_{n})$ be an exhaustion of a non compact connected surface $S$. Then every ideal boundary component of $S$ is equivalent to another of the form $(W_{n})$, where $W_{n}$ is a connected component of $S\setminus F_{n}$.
The proof is not difficult. It is a matter of showing that any ideal boundary component of $S$ is equivalent to another, where the sets are components of the complement of the surfaces of the exhaustion.
We would like to emphasise the following simple consequences of Proposition 2.5.
2.6 Corollary.
Every ideal boundary point of $b\in b(S)$ can be represented by an IBC $(W_{n})$, where $fr_{S}{W_{n}}$ is a circle.
2.7 Corollary.
If for every $n$ the number of connected components of $S\setminus F_{n}$ is less than or equal $k$ then $S$ has at most $k$ ideal boundary points.
2.1.2. Kerékjártó’s Theorem
This is a result that gives necessary and sufficient conditions for connected boundaryless surfaces to be homeomorphic.
We say that $S$ has finite genus if there exists a compact bordered surface $K$ contained in $S$ such that $S\setminus K$ is homeomorphic to a subset of the plane. In this case the genus of $S$ is defined to be the genus of $K$. Otherwise we say that $S$ has infinite genus. The genus of a connected surface may also be defined as is the maximum number of disjoint simple closed curves that can be embedded in the surface without disconnecting it.
Now we are going to define four types of orientability for a non compact surface. Assume that $S$ is not orientable. $S$ is finitely non orientable if there exists a compact bordered surface $K$ contained in $S$ such that $S\setminus K$ is orientable. Otherwise we say that $S$ is infinitely non orientable. Every compact non orientable surface is the connected sum of a compact orientable surface and one or two projective planes. If $S$ is finitely non orientable and $S\setminus K$ is orientable, we say that $S$ is of odd or even non orientability type according to $K$ is the connected sum of a compact orientable surface and one or two projective planes, respectively.
Let $b=(P_{n})\in b(S)$.
We say that $b$ is planar if some $P_{n}$ is homeomorphic to a subset of the plane. Otherwise, $b$ is called non planar, and the set of non planar ideal boundary points of $S$ is denoted by $b^{\prime}(S)$. If $b\in b^{\prime}(S)$ then the genus of $P_{n}$ is infinite for every $n$.
We say that $b$ is orientable if some $P_{n}$ is orientable. Otherwise, $b$ is called non orientable, and the set of non orientable ideal boundary points is denoted by $b^{\prime\prime}(S)$. If $b\in b^{\prime\prime}(S)$ then $P_{n}$ is infinitely non orientable for every $n$.
Obviously $b(S)\supset b^{\prime}(S)\supset b^{\prime\prime}(S)$.
2.8 Proposition.
(Kerékjártó) Let $S_{1}$ and $S_{2}$ be two non compact connected surfaces which have the same genus and the same orientability type. Then $S_{1}$ and $S_{2}$ are homeomorphic if and only if there exists a homeomorphism of $b(S_{1})$ onto $b(S_{2})$, such that $b^{\prime}(S_{1})$ and $b^{\prime\prime}(S_{1})$ are mapped onto $b^{\prime}(S_{2})$ and $b^{\prime\prime}(S_{2})$, respectively.
For a proof see Kerékjártó [10] and Theorem 1 of [15].
2.1.3. A model for non compact connected boundaryless surfaces
Let $S^{2}=\mathbb{R}^{2}\cup\{\infty\}$ be the one point compactification of the plane. We consider the Cantor ternary set $C$ as the subset of $S^{2}$ consisting of all points $(x,0)$ such that $x$ has one ternary expansion which contains no $1$’s.
Let $X\supset Y\supset Z$ be subsets of $C$ homeomorphic to $b(S)$, $b^{\prime}(S)$ and $b^{\prime\prime}(S)$, respectively (it is possible that $S$ has finite genus which is equivalent to $b^{\prime}(S)=b^{\prime\prime}(S)=\varnothing$).
If we look at $C$ as obtained by the process of removing middle thirds, then $C=\cap_{n\geq 1}J_{n}$ where $(J_{n})$ is a nested sequence of sets consisting of the union of pairwise disjoint closed intervals $I_{n}^{k}$ of length $\frac{1}{3^{n}}$, $1\leq k\leq 2^{n}$. For each $n$ choose a collection of $2^{n}$ pairwise disjoint open balls $B_{n}^{k}$ such that $I_{n}^{k}\subset B_{n}^{k}$, the centers of $B_{n}^{k}$ and $I_{n}^{k}$ coincide, and the balls have the same radius. We also want that $B_{n+1}^{l}\cap B_{n}^{k}=\varnothing$ or $B_{n+1}^{l}\subset B_{n}^{k}$. We still have that $C=\cap_{n\geq 1}\cup_{1\leq k\leq 2^{n}}B_{n}^{k}$ and for each $x\in C$ there exists a unique sequence $k_{n}$ such that $\cap_{n\geq 1}B_{n}^{k_{n}}=\{x\}$.
Every ball $B_{n}^{k}$ contains exactly two balls $B_{n+1}^{l}$ and $B_{n+1}^{l+1}$.
If $B_{n}^{k}$ contains a point of $Z$ then we choose a closed disk $D$ contained in $B_{n}^{k}$ and disjoint from the closures of $B_{n+1}^{l}$ and $B_{n+1}^{l+1}$, and make the connected sum of $S^{2}$ and a projective plane along the boundary of $D$.
If $B_{n}^{k}$ contains a point of $Y-Z$ then we choose a closed disk $D$ contained in $B_{n}^{k}$ and disjoint from the closures of $B_{n+1}^{l}$ and $B_{n+1}^{l+1}$, and make the connected sum of $S^{2}$ and a torus along the boundary of $D$.
Let $M_{0}$ be the set obtained after these connected sums and $M_{1}=M_{0}\setminus X$.
In $M_{0}$, we have that points of $Z$ are accumulated by projective planes, points of $Y\setminus Z$ are accumulated by tori and points of $M_{0}\setminus Y$ (which includes $X\setminus Y$) have neighborhoods homeomorphic to open balls. Therefore $M_{1}$ is a surface.
For each $x\in X$ there exists a unique sequence $k_{n}$ such that $\cap_{n\geq 1}B_{n}^{k_{n}}=\{x\}$. This gives a one to one correspondence between points of $x\in X$ and all ideal boundary components $(B_{n}^{k_{n}})$ of $b(M_{1})$ It is not difficult to show that this correspondence gives a homeomorphism from $b(M_{1})$ onto $X$ that takes $b^{\prime}(M_{1})$ to $Y$ and $b^{\prime\prime}(M_{1})$ to $Z$ (the final part of the proof of Theorem 2 of [15] contains a detailed proof of this fact).
From the beginning we can assume $S$ and $M_{1}$ have the same genus and orientability type. It follows from Proposition 2.8 that $S$ and $M_{1}$ are homeomorphic. To summarize, we have the following:
2.9 Proposition.
Let $S$ be a non compact connected surface of infinite genus. Then $S$ is homeomorphic to a surface obtained from the sphere $S^{2}$ by removing a totally disconnected compact subset $X$, taking a collection of pairwise disjoint closed disks $(D_{n})$ and making the connected sum of $S^{2}$ and a torus or a projective plane along the boundaries of $D_{n}$. Given a neighborhood $W$ of $X$ in $S^{2}$ all but finite many disks $D_{n}$ are contained in $W$.
See Theorem 3 of [15] and the discussion preceding it. Although $B(S)$ is not always a surface, the above description gives a very good description of what it looks like. $B(S)$ is locally homeomorphic to $\mathbb{R}^{2}$ at every point not in $b^{\prime}(S)$. Points of $b^{\prime\prime}(S)$ are accumulated by projective planes and points of $b^{\prime}(S)\setminus b^{\prime\prime}(S)$ are accumulated by tori.
Many results become very clear from Proposition 2.9. For example.
2.10 Proposition.
The following conditions about a connected boundaryless surface are equivalent.
  1. (1)
    $B(S)$ is a compact surface.
  2. (2)
    $S$ has finite genus.
  3. (3)
    Every $b\in b(S)$ is planar.
2.2. The impression of an ideal boundary point
In this subsection $S$ will be a compact connected surface.
We will consider domains $U$ of $S$ and start to investigate the relation between $b(U)$ and $fr_{S}U$.
Firstly, a simple consequence of Proposition 2.10.
2.11 Proposition.
Let $S$ be a compact connected surface and $U$ a domain of $S$. Then every ideal boundary point of $U$ is planar and $B(U)$ is a compact surface.
Let $b=(P_{n})\in b(U)$. The impression of $b$ in $S$ is the set $Z(b):=\cap_{n}cl_{S}{P_{n}}$.
2.12 Proposition.
Let $b=(P_{n})\in b(U)$. $Z(b)$ is the set of points $x\in S$ such that there exists a sequence $(x_{n})$ in $U$ such that $x_{n}\rightarrow b\in b(U)$ in $B(U)$ and $x_{n}\rightarrow x$ in $S$.
Proof.
Let $x\in S\,|\,\exists\,(x_{n})\text{ in }U\text{ such that }x_{n}\rightarrow b\in b(U)\text{ in }B(U)\text{ and }x_{n}\rightarrow x\text{ in }S$. For every $k$ there exists $n_{k}$ such that $x_{n}\in P_{k}$ for $n\geq n_{k}$. Since $x_{n}\rightarrow x\text{ in }S$ we have that $x\in cl_{S}P_{k}$ for every $k.$
Conversely, assume that $x\in cl_{S}P_{n}$ for every $n$. Let $(W_{k})$ be a fundamental system of neighborhoods of $x$ in $S$. For each $k$ take a point $x_{k}\in W_{k}\cap P_{k}.$ Obviously $x_{k}\rightarrow x$ in $S$. Since $x_{k}\in P_{k}$ for all $k$ we have that $x_{k}\rightarrow b$ in $U$. ∎
Now we would like to prove the following.
2.13 Proposition.
Let $S$ be a compact connected surface and $U$ a domain of $S$. Then $fr_{S}U=\cup_{b\in b(U)}Z(b)$.
Proof.
Let $x\in Z(b)$, where $b=(P_{n})\in b(U)$.
Since $P_{n}\subset U$ for every $n$, we have that $Z(b)\subset cl_{S}U$. Assume by contradiction that $x\in U$. Then $x\in U\cap cl_{S}P_{n}=cl_{U}P_{n}$ for every $n$, contradicting the fact that $(P_{n})$ is an IBC of $U$. This proves that $x\in fr_{S}U$.
Conversely, suppose that $x\in fr_{S}U$.
Let $(x_{k})$ be a sequence in $U$ such that $x_{k}\rightarrow x$ in $S$ and consider an exhaustion $(F_{n})$ of $U$. Recall that $U\setminus F_{n}$ has a finite number of components for every $n$. This implies that, for each $n$, at least one of the components of $U\setminus F_{n}$ contains an infinite number of elements of $(x_{k})$.
Claim. For every $m$, there exists open connected sets $P_{1}\supset P_{2}\supset\dots\supset P_{m}$ such that:
  1. (1)
    $P_{j}$ is a component of $U\setminus F_{j}$, for $1\leq j\leq m$.
  2. (2)
    $\{k\mid x_{k}\in P_{m}\}$ is infinite.
The claim is obvious for $m=1$. Assume that the claim is true for $1\leq j\leq m$. Then at least one of the components of $U\setminus F_{m+1}$ contained in $P_{m}$, say $P_{m+1}$, has the property that $\{k\mid x_{k}\in P_{m+1}\}$ contains an infinite subset of $\{k\mid x_{k}\in P_{m}\}$. This proves the claim.
From the claim we conclude that $(P_{n})$ is an IBC of $U$ that defines an ideal boundary point $b\in b(U)$. We also conclude that there exists an increasing sequence $k_{n}\rightarrow\infty$ such that $x_{k_{n}}\in P_{n}$ for every $n$. We have that $x_{k_{n}}\rightarrow b$, and since $x_{k_{n}}\rightarrow x$ in $S$ we have that $x\in Z(b)$. ∎
2.3. Domains whose frontier has a finite number of connected components
In this subsection $S$ will be a compact surface. We would like to prove the following result.
2.14 Theorem.
Let $S$ be a compact connected surface and $U$ be a domain of $S$. Then the following conditions on $U$ are equivalent:
  1. (1)
    $U$ is a residual domain of a compact set $K$ such that $fr_{S}K$ has a finite number of connected components.
  2. (2)
    $U$ has a finite number of ideal boundary points.
  3. (3)
    $fr_{S}{U}$ has a finite number of connected components.
These conditions are not equivalent if $S$ is not a surface and could be false in more general compact connected topological spaces.
Consider the following compact connected subset of $\mathbb{R}^{2}$: $X=([0,1]\times\{0,2\})\cup(C\times[0,2])$, where $C$ is an infinite closed subset of $[0,1]$. If $K=\{(x,y)\in X\thinspace|\thinspace y\leq 1\}$ then $K$ is compact and connected. $K$ has only one residual domain $U$, but $fr_{S}{U}$ has an infinite number of components and ends.
2.4. An upper bound on the number of ideal boundary points of a domain
Now we would like to give an upper bound on the number of ideal boundary points of a domain $U$ contained in a compact connected surface $S$.
Our result can be used to give an elementary proof of Lemma 2.3 of [13], where no upper bound is explicitly given.
2.15 Theorem.
Let $K$ be a compact subset of a compact connected surface $S$ and $U$ a residual domain of $K$. If $K$ has $m$ connected components then $U$ has at most $m(g+1)$ ideal boundary points, where $g$ is the genus of $S$.
Before the proof, we would like to note that the upper bound is sharp in the sense that, on any compact connected surface, there exist examples where it is reached. In a compact connected orientable surface $S$ of genus $g$, take one meridian $\gamma_{i}$ around each handle of $S$ and connected them with a simple curve $\alpha$ in the simplest way. Then $K=\alpha\cup\gamma_{1}\cup\dots\cup\gamma_{g}$ is a compact connected subset of $S$ and $U=S\setminus K$ is connected. It easy to see that $U$ has $g+1$ ideal boundary points. In the non orientable case we can give a similar example by thinking of $S$ as the sphere with $g$ Mobius strips attached to it, and taking the curves $\gamma_{i}$ as the ”the central circles” of the strips.
We are going to divide the proof into a few lemmas.
2.16 Lemma.
We may assume that $S=K\sqcup U$, where $U$ is the only residual domain of $K$.
Proof.
Let $W=\sqcup_{\lambda\in L}V_{\lambda}$ be the union of the components of $S\setminus K$ different from $U$ and $K^{\prime}=S\setminus U=K\sqcup W$. Obviously, $S=K^{\prime}\sqcup U$.
Since $K\subset K^{\prime}$ every component of $K$ is contained in some component of $K^{\prime}$.
Claim. Every component of $K^{\prime}$ contains at least one component of $K$.
To prove the claim it is enough to show that every component $C$ of $K^{\prime}$ intersects $K$. Assume by contradiction that $C\cap K=\varnothing$. Then $C\subset W$. The components of $W$ are the sets $(V_{\lambda})_{\lambda\in L}$. Therefore $C\subset V_{\lambda_{0}}$ for some $\lambda_{0}\in L$. But $C$ is a maximal connected subset of $K^{\prime}$, and therefore $C=V_{\lambda_{0}}$. This implies that $C$ is open in $S$.
On the other hand, being a component, $C$ is closed in $K^{\prime}$ and $C$ is closed in $S$. This contradicts that $S$ is connected, which proves the claim.
From the claim we conclude that the number of components of $K^{\prime}$ is less than or equal to the number of components of $K$. Therefore we could use the decomposition $S=K^{\prime}\sqcup U$ to prove Theorem 2.15. ∎
Now we would like to calculate an upper bound for the number of ideal boundary points of $U$. Let $(F_{n})_{n\geq 1}$ be an exhaustion of $U$.
2.17 Lemma.
If $W$ is a component of $S\setminus F_{n}$ then $W$ contains at least one component of $K$.
Proof.
Since $K\subset S\setminus F_{n}$, we have that every component of $K$ is contained in some component of $S\setminus F_{n}$. It is enough to show that $W\cap K\neq\varnothing$.
Let us assume by contradiction that $W\cap K=\varnothing$.
It follows that $W\subset U$ and that $W$ is contained in a connected component $V$ of $U\setminus F_{n}$. But every component of $U\setminus F_{n}$ is contained in some component of $S\setminus F_{n}$. From this we conclude that $W=V$.
We have that $cl_{S}W=W\cup\sigma_{1}\cup...\cup\sigma_{l}$, where each $\sigma_{i}$ is a circle of $\partial{F_{n}}\subset U$. From this we conclude that $cl_{S}W\subset U$ and that $cl_{U}V=cl_{U}W=U\cap cl_{S}W=cl_{S}W$ is compact. But this contradicts the fact that $(F_{n})$ is an exhaustion of $U$.
Therefore $W$cap $K$neq $\emptyset$ and every connected component of $S\setminus F_{n}$ contains at least one component of $K$. ∎
Let $E_{1},...,E_{k}$ be the components of $S\setminus F_{n}$. From Lemma 2.17 we have that $k\leq m$.
For $1\leq i\leq k$, we have that each $E_{i}$ is an open set whose frontier in $S$ is a union of circles of $\partial F_{n}$. We write $fr_{S}E_{i}=C_{i1}\sqcup\dots\sqcup C_{i\nu_{i}}$, where $\nu_{i}\geq 1$.
Since $\partial F_{n}=fr_{S}F_{n}=fr_{S}E_{1}\sqcup\dots\sqcup fr_{S}E_{k}$, we have that the number of components of $\partial F_{n}$ is $\nu_{1}+\dots+\nu_{k}$.
With this notation we have that
$S=int_{S}F_{n}\sqcup fr_{S}F_{n}\sqcup(E_{1}\sqcup\dots\sqcup E_{k})$
or
$S=int_{S}F_{n}\sqcup\big{(}\sqcup_{1\leq i\leq k}(E_{i}\sqcup fr_{S}E_{i})\big{)}$.
2.18 Lemma.
For each $i$, $1\leq i\leq k$, and any subset $\{j_{1},\dots,j_{s}\}$ of $\{1,\ldots,\nu_{i}\}$ we have that the set $int_{S}F_{n}\sqcup E_{i}\sqcup\big{(}C_{ij_{1}}\sqcup\dots\sqcup C_{ij_{s}}\big{)}$ is connected.
Proof.
We have that $E_{i}$ is connected and $E_{i}\subset E_{i}\sqcup\big{(}C_{ij_{1}}\sqcup\dots\sqcup C_{ij_{s}}\big{)}\subset cl_{S}E_{i}$. Therefore $E_{i}\sqcup\left(C_{ij_{1}}\sqcup\dots\sqcup C_{ij_{s}}\right)$is connected.
Similarly, $int_{S}F_{n}$ is connected, and since
$int_{S}F_{n}\subset int_{S}F_{n}\sqcup(C_{ij_{1}}\sqcup\dots\sqcup C_{ij_{s}})\subset cl_{S}F_{n}$
we have that $int_{S}F_{n}\sqcup(C_{ij_{1}}\sqcup\dots\sqcup C_{ij_{s}})$ is connected. ∎
We claim that $\nu_{i}\leq g+1$, for $1\leq i\leq k$.
If we had $\nu_{l}\geq g+2$ for some $l\in\{1,\dots,k\}$, then $S\setminus(C_{l1}\sqcup\dots\sqcup C_{l(g+1)})$ would be disconnected. But $S\setminus(C_{l1}\sqcup\ldots\sqcup C_{l(g+1)})=$
$\big{[}\bigsqcup_{1\leq i\leq k,i\neq l}\big{(}int_{S}F_{n}\sqcup(E_{i}\sqcup fr_{S}E_{i})\big{)}\big{]}\bigsqcup\big{[}int_{S}F_{n}\sqcup(E_{l}\sqcup C_{l(g+2)}\sqcup\dots\sqcup C_{l\nu_{l}})\big{]}$
and the second side of the equality is a connected set by Lemma 2.18, a contradiction.
The number of components of $\partial F_{n}$ is $\eta_{n}:=\nu_{1}+...+\nu_{k}$. It follows that $\eta_{n}=\nu_{1}+...+\nu_{k}\leq k(g+1)\leq m(g+1)$.
From Corollary 2.7 we have that $U$ has at most $m(g+1)$ ideal boundary points.
2.5. Proof of Theorem 2.14: equivalent conditions on the frontier of a domain
Let $S$ be a compact connected surface and $U$ be a domain of $S$. We would like to prove that the following conditions on $U$ are equivalent:
  1. (1)
    $U$ is a residual domain of a compact set $K$ such that $fr_{S}K$ has a finite number of connected components.
  2. (2)
    $U$ has a finite number of ideal boundary points.
  3. (3)
    $fr_{S}{U}$ has a finite number of connected components.
$(1)$ implies $(2)$.
If $U$ is a residual domain of a compact set $K$ such that $fr_{S}K$ has a finite number of connected components, then by Theorem 2.15 we have that $b(U)$ is finite.
$(2)$ implies $(3)$.
From Proposition 2.13 we know that $fr_{S}U=\cup_{b\in b(U)}Z(b)$. Since $b(U)$ is finite and each set $Z(b)$ is connected, we have that $fr_{S}U$ has a finite number of components.
$(3)$ implies $(1)$. If we take $K$ as $fr_{S}U$, the result follows from the simple fact that $U$ is always a residual domain of $fr_{S}U$.
2.6. The prime ends compactification
In this subsection we will consider a compact connected surface $S$ and a domain $U\subset S$ such that $fr_{S}U$ has a finite number of connected components. We will describe the prime ends compactification of $U$.
As we saw in Theorem 2.14, $fr_{S}U$ has finite number of connected components if and only if $b(U)$ is finite. Observe that $U$ is a surface finite genus and by Proposition 2.10 $B(U)$ is a compact surface. Therefore we are in the setting of Mather’s work [14] and we will make a brief presentation of his theory of prime ends. The basis of Mather’s work are [4] and [5].
The basic idea is the following. If $b$ is an ideal boundary point of $U$ and $Z(b)$ has more that one point, we call $b$ regular. Otherwise the $b$ is just a puncture of $U$ in $S$. For each regular ideal boundary point $b\in b(U)$, we replace the impression $Z(b)$ of $b$ by a circle $C(b)$, and obtain a compact connected surface with boundary.
The reader can find the proof of all results that we are going to list in this subsection in [14].
A chain is a sequence $(V_{i})_{i\geq 1}$ of open connected subsets of $U$ such that
  1. (1)
    $V_{i}\supset V_{i+1}$ for every $i$.
  2. (2)
    $fr_{U}V_{i}$ is nonempty and connected for every $i$.
  3. (3)
    $cl_{S}(fr_{U}V_{i})\cap cl_{S}(fr_{U}V_{j})=\varnothing$ for $i\neq j$.
A chain $(W_{j})$ divides $(V_{i})$ if for every $i$ there exists $j$ such that $W_{j}\subset V_{i}$. Two chains are equivalent if each divides the other. A chain is prime if any chain which divides it is equivalent to it. A sufficient condition for a chain $(V_{i})$ to be prime is that there exists $p\in S$ such that $fr_{U}V_{i}\rightarrow p$ (every neighborhood of $p$ contains all but finitely many of the sets $fr_{U}V_{i}$).
A prime point is an equivalence class of prime chains.
Let $x\in U$ and consider a family of closed disks $D_{1}\supset D_{2}\supset\dots$ in $U$ such that $D_{i+1}\subset int_{U}D_{i}$ and $\cap D_{i}=\{x\}$. Then $(int_{U}D_{i})$ is a prime chain that defines a prime point denoted by $\omega(x)$.
A prime end is a prime point which is not of the form $\omega(x)$ for any $x\in U$. The set of prime points of $U$ is denoted by $E(U)$.
We consider a topology on $E(U)$ defined as follows. Let $V$ be an open subset of $U$ and denote by $V^{\prime}$ the set of prime points of $U$ whose representing chains are eventually contained in $V$. The collection of sets $V^{\prime}\subset E(U)$ such that $V$ is an open subset of $U$ is a basis of a topology on $E(U)$.
The function $\omega:U\rightarrow E(U)$ is a homeomorphism from $U$ onto an open subset of $E(U)$, and we simply identify $U$ with $\omega(U)$.
For $e\in E(U)$ let $\alpha(e)=\cap_{i}cl_{B(U)}V_{i}$, where $(V_{i})$ is a chain representing $e$. Then $\alpha(e)$ consists of one point and $\alpha:E(U)\rightarrow B(U)$ is a continuous function whose restriction to $U$ is the inclusion (or more precisely, $\alpha\circ\omega$ is the identity on $U$).
$E(U)$ is a compact connected surface with boundary.
We have that $\partial E(U)=\cup_{b\in b_{reg}(U)}\alpha^{-1}(b)$. When $b\in b_{reg(U)}$, we have that $\alpha^{-1}(b)$ is homeomorphic to a circle, called the Caratheodory circle of prime ends associated with $b$. We also denote it by $C(b)$.
We have a decomposition $E(U)=U\sqcup\alpha^{-1}\big{(}b(U)\big{)}$, where $\alpha^{-1}\big{(}b(U)\big{)}$ is the set of all prime ends of $U$. If $b\in b(U)$ is not regular, then the prime end $\alpha^{-1}(b)$ is just a point of $E(U)\setminus\partial E(U)$.
Let $e\in E(U)$ and $(V_{i})$ be a chain representing $e$. The set $Y(e)=\cap_{i}cl_{S}(V_{i})$ is called the impression of $e$. The definition does not depend on the representing chain, and $Y(e)$ is a compact, connected, non-empty subset of $S$.
We say that $p\in S$ is a principal point of $e$ if there is a chain $(V_{i})$ which represents $e$ such that $fr_{U}V_{i}\rightarrow p$. The set of principal points of $e$ is called the principal set of $e$ and is denoted by $X(e)$. It is a non-empty, compact, connected subset of $S$, and $X(e)\subset Y(e)\subset Z(b)$, where $b=\alpha(e)$.
We will need a result that is probably known by experts, but as far as we know is not written anywhere.
2.19 Proposition.
Let $S$ be a compact connected surface and $U\subset S$ a domain such that $fr_{S}U$ has finite number of connected components. If $b$ is a regular ideal boundary point of $U$ then $Z(b)=\cup_{e\in C(b)}Y(e)$.
Proof.
We already know that $Y(e)\subset Z(b)$ for every $e\in C(b)$.
Let’s prove that $Z(b)\subset\cup_{e\in C(b)}Y(e)$.
Let $x\in Z(b)$ where $b=(P_{n})$. We know that $x\in Z(b)$ if and only if there exits a sequence $(x_{k})$ in $U$, such that $x_{k}\rightarrow x$ in $S$ and $x_{k}\rightarrow b\in b(U)$ in $B(U)$.
This last convergence means that
(1) for every $\text{for every }n\text{ there exists }k_{n}\text{ such that }k\geq k_{n}\text{ implies }x_{k}\in P_{n}.$
From this we conclude that $x\notin U$.
We have that $(x_{k})$ is a sequence in $E(U)=U\sqcup\alpha^{-1}\big{(}b(U)\big{)}$. Taking a subsequence if necessary, we may assume that $x_{k}\rightarrow e\in E(U)$. We can not have $e\in U$, otherwise $e=x$ and $x\notin U$. Therefore $e\in\alpha^{-1}\big{(}b(U)\big{)}$, which is the set of all prime ends of $U$.
By the construction of $E(U)$, for every $n$ we have that $C(b)\sqcup P_{n}$ is a neighborhood of $C(b)$ in $E(U)$. We can assume that this neighborhood contains no other prime ends than those of $C(b)$.
Therefore by (1) we have that $e\in C(b)$.
If $(V_{i})$ is a chain representing $e$, then for every $i$ there exists $k_{i}$ such that $x_{k}\in V_{i}\subset cl_{S}V_{i}$ for $k\geq k_{i}$. From this we conclude that $x\in cl_{S}V_{i}$ for every $i$ and $x\in Y(e)=\cap_{i}cl_{S}V_{i}$, where $e\in C(b)$. ∎
Let $\beta:(0,1]\rightarrow U$ be a path such that $\underset{t\rightarrow 0}{\lim}\thinspace\beta(t)=p$ in $S$ for some $p\in fr_{S}U$. Then there exist $e\in E(U)$ and $b\in b(U)$ such that $\underset{t\rightarrow 0}{\lim}\thinspace\beta(t)=e$ in $E(U)$ and $\underset{t\rightarrow 0}{\lim}\thinspace\beta(t)=b$ in $B(U)$.
If $X(e)=\{p\}$ for some $p\in fr_{S}U$, we say that $e$ is an accessible prime end. We have that $e$ is an accessible prime end if and only if there exists a path $\beta:(0,1]\rightarrow U$ such that $\underset{t\rightarrow 0}{\lim}\beta(t)=p$ in $S$ and $\underset{t\rightarrow 0}{\lim}\beta(t)=e$ in $E(U)$.
2.7. A preparation lemma: chains whose frontiers are arcs of circles
We would like to show that accessible prime ends can be represented by chains whose frontiers are contained in arcs of circles.
[…]
As before $S$ will be a compact connected surface, $U\subset S$ will be a domain whose frontier contains a finite number of components and $E(U)$ its prime ends compactification. Let $e$ be an accessible prime end of $U$ and $p$ its principal point. In a neighborhood of $p$ we consider continuous coordinates with $p$ at the origin.
In this subsection we will prove the following result.
2.20 Lemma.
Let $S$ be a compact connected surface, $U\subset S$ a domain whose frontier contains a finite number of components. Let $e$ be an accessible prime end of $U$ and $p$ its principal point. Then we have the following:
Then there exists $\delta>0$ such that for any decreasing sequence $(r_{n})_{n\geq 1}$ contained in $(0,\delta)$ with $\underset{n\rightarrow\infty}{\lim}r_{n}=0$, there exists a chain $(V_{n})$ representing $e$ such that $fr_{U}V_{n}\subset C_{r_{n}}$.
We start by taking a path $\beta:(0,1]\rightarrow U$ such that $lim_{t\rightarrow 0}\beta(t)=p$ in $S$ and $lim_{t\rightarrow 0}\beta(t)=e$ in $E(U)$.
There exits $b\in b(U)$ such that $lim_{t\rightarrow 0}\beta(t)=b$ in $B(U)$.
This follows from that fact that the limit set of $\beta$ as $t\rightarrow 0$ in $B(U)$, $\cap_{\epsilon>0}cl_{B(U)}\beta(0,\epsilon)$, is a non empty compact connected subset of $B(U)$. Since $p\in fr_{S}U$, we have that $\cap_{\epsilon>0}cl_{B(U)}\beta(0,\epsilon)$ does not intersect $U$, and therefore is contained in $b(U)=B(U)\setminus U$. On the other hand $b(U)$ is finite, and from this we conclude that the limit set of $\beta$ as $t\rightarrow 0$ in $B(U)$ is a single point $b\in b(U)$.
There exist a closed disk $D$ in $B(U)$ such that $b\in D^{\circ}$, $D-\{b\}\subset U$ and $\beta(0,1]\cap\partial D\neq\varnothing$.
Let $t_{0}=\inf\{t\in(0,1]$ |||| $\beta(t)\in\partial D\}$. It follows that $\beta(t_{0})\in\partial D$ and $\beta(0,t_{0})\subset D^{\circ}$.
Since $\partial D$ is a compact subset of $U$ and $p\in fr_{S}U$, there exists $\delta>0$ such that $B_{r}\cap\partial D=\varnothing$ if $r<\delta$.
2.21 Proposition.
For any $r\in(0,\delta)$, let $c$ and $d$ such that $0<c<d\leq t_{0}$ and assume that $\beta(c)\in B_{r}^{\circ}$ and $\beta(d)\notin B_{r}$.
Then there exists a connected component $\xi$ of $C_{r}\cap U$ that separates $U$ into two open connected subsets $V$ and $V^{\prime}$, such that $\beta(c)\in V$, $\beta(d)\in V^{\prime}$ and $fr_{U}{V}=fr_{U}{V^{\prime}}=\xi$.
Proof.
Note that the components of $C_{r}\cap U$ form an open cover of $\beta[c,d]\cap C_{r}$, which is compact. Therefore $\beta[c,d]$ intersects only finitely many components of $C_{r}\cap U$, which we denote by $\lambda_{1},...,\lambda_{n}$.
For $i=1,...,n$ we have that $\beta[c,d]\subset D-\{b\}$ and therefore $\lambda_{i}\cap(D-\{b\})\neq\varnothing$. On the other hand, since $\lambda_{i}\subset C_{r_{i}}$ we have that $\lambda_{i}\cap\partial D=\varnothing$. Therefore $\lambda_{i}\subset D^{\circ}-\{b\}$.
We have that $\lambda_{i}$ is an open arc in $B(U)$ with end points at $b$ and therefore $\lambda_{i}\cup\{b\}\subset D$ is homeomorphic to a circle. From this we conclude that $U-\lambda_{i}$ has two components for $1\leq i\leq n$.
Let $\lambda_{0}$ be the union of the components of $C_{r}\cap U$ different from $\lambda_{1},...,\lambda_{n}$.
We are going to assume that $\beta(c)$ and $\beta(d)$ belong to the same component of $U-\lambda_{i}$ for $1\leq i\leq n$ and obtain a contradiction.
We are going to prove that for every $i=0,1,...,n$, there exists a path $\alpha_{i}:[c,d]\rightarrow U-(\lambda_{0}\cup...\cup\lambda_{i})$ from $\beta(c)$ to $\beta(d)$.
If we take $\alpha_{0}:[c,d]\rightarrow U-\lambda_{0}$ defined by $\alpha_{0}(t)=\beta(t)$, then $\alpha_{0}(c)=\beta(c)$, $\alpha_{0}(d)=\beta(d)$, and the statement is true for $i=0$.
Assume the statement is true for $k=0,...,i$ where $i<n$.
Then there exists a path $\alpha_{i}:[c,d]\rightarrow U-(\lambda_{0}\cup...\cup\lambda_{i})$ from $\beta(c)$ to $\beta(d)$. We would like to construct a path $\alpha_{i+1}:[c,d]\rightarrow U-(\lambda_{0}\cup...\cup\lambda_{i}\cup\lambda_{i+1})$ from $\beta(c)$ to $\beta(d)$.
If $\alpha_{i}$ does not intersect $\lambda_{i+1}$ then we just take $\alpha_{i+1}=\alpha_{i}$. Now suppose that $\alpha_{i}\cap\lambda_{i+1}\neq\varnothing$.
Let $u=\inf\{t\thinspace|\thickspace\alpha_{i}(t)\in\lambda_{i+1}\}$ and $v=\sup\{t\thinspace|\thickspace\alpha_{i}(t)\in\lambda_{i+1}\}$. Recall that we are assuming that $\beta(c)$ and $\beta(d)$ belong to the same component of $U-\lambda_{i+1}$. Therefore, if $u^{\prime}<u$ and $v^{\prime}>v$ then $\alpha[c,u^{\prime}]$ and $\alpha[v^{\prime},d]$ are contained in the same component of $U-\lambda_{i+1}$.
Hence we can choose $u^{\prime}$ and $v^{\prime}$ close enough to $u$ and $v$ satisfying the following:
  • $\alpha(u^{\prime})$ and $\alpha(v^{\prime})$ are close to $\lambda_{i+1}$ and on the same side of it.
  • $\alpha(u^{\prime})$ and $\alpha(v^{\prime})$ are the end points of an arc $\eta$ contained in $C_{r^{\prime}}\cap U$ with $r^{\prime}\neq r$.
It follows that the arc $\eta$ does not intersect $\lambda_{0}\cup...\cup\lambda_{i}\cup\lambda_{i+1}$.
If we define $\alpha_{i+1}$ by replacing the part of $\alpha_{i}$ between $\alpha_{i}(u^{\prime})$ and $\alpha_{i}(v^{\prime})$ by $\eta$ we obtain the desired path $\alpha_{i+1}:[c,d]\rightarrow U-(\lambda_{0}\cup...\cup\lambda_{i}\cup\lambda_{i+1})$ from $\beta(c)$ to $\beta(d)$. Therefore the statement is true for $k=i+1$.
Hence there exists a path $\alpha_{n}:[c,d]\rightarrow U-(\lambda_{0}\cup\lambda_{1}\cup...\cup\lambda_{n})$ from $\beta(c)$ to $\beta(d)$.
This is a contradiction, since any path in $U$ from $\beta(c)$ to $\beta(d)$ must intersect $C_{r}\cap U=\lambda_{0}\cup\lambda_{1}\cup...\cup\lambda_{n}$. ∎
Returning to the proof of the preparation Lemma 2.20, let $(r_{n})_{n\geq 1}$ contained in $(0,\delta)$ be any decreasing sequence such that $\underset{n\rightarrow\infty}{\lim}r_{n}=0$.
For $i\geq 1$, let $t_{i}=\inf\{t\in(0,1]\thinspace|\thinspace\beta(t)\in C_{r_{i}}\}$. Then $\beta(t_{i})\in C_{r_{i}}$ and $\beta(0,t_{i})\subset B_{r_{i}}^{\circ}$.
Now we would like to construct the chain.
$V_{1}$ is defined in the following way. We have that $\beta(t_{0})\notin B_{r_{1}}$ and $\beta(0,t_{1})\subset B_{r_{1}}^{\circ}$. By Proposition 2.21, there exists a component $\xi_{1}$ of $C_{r_{1}}\cap U$ that separates $U$ into two components $V_{1}$ and $V^{\prime}_{1}$, such that $\beta(0,t_{1})\subset V_{1}$, $\beta(t_{0})\notin B_{r_{1}}$ and $fr_{U}{V_{1}}=\xi_{1}$. Note that $\beta(t_{0})\notin cl_{U}V_{1}$.
Suppose now that we have defined $V_{1}\supset V_{2}\supset...\supset V_{n}$ with the following properties for $1\leq i\leq n$:
  1. (1)
    $\beta(0,t_{i})\subset V_{i}$.
  2. (2)
    $fr_{U}V_{i}=\xi_{i}$ is a connected component of $C_{r_{i}}\cap U$.
  3. (3)
    $\beta(t_{i-1})\notin cl_{U}V_{i}$.
$V_{n+1}$ is defined in the following way.
We have that $\beta(t_{n})\notin B_{r_{n+1}}$ and $\beta(0,t_{n+1})\subset B_{r_{n+1}}^{\circ}$. By Proposition 2.21, there exists a component $\xi_{n+1}$ of $C_{r_{n+1}}\cap U$ that separates $U$ into two components $V_{n+1}$ and $V^{\prime}_{n+1}$, such that $\beta(0,t_{n+1})\subset V_{n+1}$, $\beta(t_{n})\notin B_{r_{n+1}}$ and $fr_{U}{V_{n+1}}=\xi_{n+1}$. Note that $\beta(t_{n})\notin cl_{U}V_{n+1}$.
Therefore $V_{n+1}$ satisfies $(1)$, $(2)$ and $(3)$ above.
We still need to prove that $V_{n+1}\subset V_{n}$. To help us with this, let’s prove a very useful Proposition.
2.22 Proposition.
Let $W_{1}$ and $W_{2}$ be open connected subsets of $U$, both non empty and different from $U$. Assume that
  1. (1)
    $W_{1}\cap W_{2}\neq\varnothing$.
  2. (2)
    $W_{1}\cap(U\setminus W_{2})\neq\varnothing$.
  3. (3)
    $W_{2}\cap(U\setminus W_{1})\neq\varnothing$.
  4. (4)
    $fr_{U}W_{1}$ and $fr_{U}W_{2}$ are connected and disjoint.
    Then we have the following:
    1. (a)
      $fr_{U}(W_{1})\subset W_{2}$ and $fr_{U}(W_{2})\subset W_{1}$.
    2. (b)
      $U=W_{1}\cup W_{2}$.
Proof.
Since $fr_{U}(W_{1})$ is connected and does not intersect $fr_{U}(W_{2})$, we must have either $fr_{U}(W_{1})\subset W_{2}$ or $fr_{U}(W_{1})\subset U\setminus W_{2}$. But $W_{2}$ is a connected set that intersects both $W_{1}$ and $U\setminus W_{1}$, and therefore we have $W_{2}\cap fr_{U}(W_{1})\neq\varnothing$. This implies that $fr_{U}(W_{1})\subset W_{2}$. Similarly $fr_{U}(W_{2})\subset W_{1}$, which proves (a).
In order to prove b), we are going to show that $W_{1}\cup W_{2}$ is open and closed in $U$. It is obviously open. Let $(a_{n})$ be a sequence in $W_{1}\cup W_{2}$ converging to a point $a$. We may assume that $a_{n}\in W_{1}$ for infinitely many values of $n$, and therefore $a\in cl_{U}(W_{1})$. If $a\in W_{1}$ we are done. If not, we have $a\in fr_{U}(W_{1})$, and from part a) $a\in W_{2}$. ∎
Now we prove that $V_{n+1}\subset V_{n}$.
We are going to apply Proposition 2.22 to $V_{n}=W_{1}$ and $V_{n+1}=W_{2}$.
We have that $\beta(0,t_{n+1})\subset V_{n}\cap V_{n+1}$, and therefore, $V_{n}$ and $V_{n+1}$ satisfy hypothesis (1) of Proposition 2.22. Since $\xi_{n}\cap\xi_{n+1}=\varnothing$ we have that hypothesis (4) of Proposition 2.22 is satisfied. Since $\beta(t_{n})\notin cl_{U}V_{n+1}$ we have $\beta(t_{n}-\delta^{\prime},t_{n})\cap V_{n+1}=\varnothing$ for $\delta^{\prime}$ small enough. Therefore $\beta(t_{n}-\delta^{\prime},t_{n})\subset V_{n}\setminus V_{n+1}$ and hypothesis (2) of Proposition 2.22 is also satisfied. We have that $\beta(t_{0})\notin cl_{U}V_{1}$ implying that $\beta(t_{0})\notin V_{n}\cup V_{n+1}$ and conclusion (b) of Proposition 2.22 is not satisfied. Therefore hypothesis (3) of Proposition 2.22 is not satisfied and $V_{n+1}\subset V_{n}$.
So the chain $(V_{n})$ is well defined.
Since $fr_{U}V_{n}\subset C_{r_{n}}$ we have that $fr_{U}V_{n}\rightarrow p$ and $(V_{n})$ is prime. Let $e^{\prime}$ be the prime end it defines. Since $\beta(0,t_{n})\subset V_{n}$ for every $n$ we have $\underset{t\rightarrow 0}{\lim}\thinspace\beta(t)=e=e^{\prime}$ in $E(U)$.
Hence $(V_{n})$ represents $e$, which proves the preparation Lemma 2.20.
2.8. Homeomorphisms and fixed points
In this subsection, $S$ will be a connected compact surface and $f:S\rightarrow S$ a homeomorphism of $S$. We also assume that $f$ preserves a finite measure $\mu$ which is positive on open sets. As before $U$ will be domain such that $fr_{S}U$ is compact and has a finite number of connected components.
Assume that $U$ is invariant under $f$.
We have that $f$ maps an ideal boundary component of $U$ to another ideal boundary component of $U$. Therefore the restriction of $f$ to $U$ extends to a homeomorphism $f_{*}:B(U)\rightarrow B(U)$ and the points of $b(U)$ are periodic under $f_{*}$.
Since $f$ maps irreducible chains to irreducible chains, the restriction of $f$ to $U$ also extends to a homeomorphism $\hat{f}:E(U)\rightarrow E(U)$ and the circles of prime ends of $U$ are all periodic under $\hat{f}$.
If $C(b)$ is the circle of prime ends associated with a regular $b\in b(U)$, then $\hat{f}\big{(}C(b)\big{)}=C\big{(}f_{*}(b)\big{)}$ and the periods of $b$ and $C(b)$ are the same. For simplicity we will assume that ideal boundary points are fixed and all circles of prime ends are invariant.
Recall that if $e\in C(b)$ then $X(e)\subset Y(e)\subset Z(b)$.
We start with the following result.
2.23 Lemma.
Let $b\in b(U)$ be a regular ideal boundary points of $U$ and assume that $\hat{f}\big{(}C(b)\big{)}=C(b)$. Suppose that $e\in C(b)$ is a fixed prime end of $U$, $p$ a principal point of $e$ and $(V_{i})$ a chain representing $e$ such that $fr_{U}(V_{i})\rightarrow p$.
There exists $i_{0}$ such that $fr_{U}(V_{i})\cap fr_{U}\big{(}f(V_{i})\big{)}\neq\varnothing$ for $i\geq i_{0}$.
Proof.
We have that $(fV_{i})$ is a chain that represents $\hat{f}(e)$. Since $e$ is a fixed point of $\hat{f}$ we have that $(fV_{i})$ and $(V_{i})$ are equivalent.
Let $i_{0}$ be such that $fV_{i}\subset V_{1}$ for $i\geq i_{0}$. We are going to assume that $fr_{U}(V_{i})\cap fr_{U}(fV_{i})=\varnothing$ for some $i\geq i_{0}$ and apply Lemma 2.22 to $V_{i}$ and $fV_{i}$ to obtain a contradiction.
So we are assuming that hypothesis $(4)$ of Lemma 2.22 is true. The equivalence of $(V_{i})$ and $(fV_{i})$ implies that $V_{i}\cap fV_{i}\neq\varnothing$ for every $i\geq 1$ and hypothesis $(1)$ is also true. We have $V_{i}\subset V_{1}$ for $i\geq i_{0}$ and therefore $V_{i}\cup fV_{i}\subset V_{1}$ for $i\geq i_{0}$. Since $V_{1}$ is strictly contained in $U$ we have that conclusion $(b)$ is false.
This implies that hypotheses $(2)$ and $(3)$ can not both hold. So we have $V_{i}\subset fV_{i}$ or $fV_{i}\subset V_{i}$. We are going to assume that $fV_{i}\subset V_{i}$ and show that $\mu(fV_{i})<\mu(V_{i})$. The other case is analogous.
Since $fr_{U}(fV_{i})$ is connected and $fr_{U}(fV_{i})\cap fr_{U}V_{i}=\varnothing$, we have that $fr_{U}(fV_{i})\subset V_{i}$ or $fr_{U}(fV_{i})\subset U-V_{i}$.
We have that $fr_{U}(fV_{i})\subset U-V_{i}$ can not happen. If we had $fr_{U}(fV_{i})\subset U-V_{i}$ then $fr_{U}(fV_{i})\subset U-cl_{U}V_{i}$. For any $x\in fr_{U}(fV_{i})$ there would exist a neighborhood $W$ of $x$ in $S$ disjoint from $cl_{U}V_{i}$. This is a contradiction since $W$ would contain points of $fV_{i}\subset V_{i}$.
Therefore $fr_{U}(fV_{i})\subset V_{i}$ implying that $cl_{U}(fV_{i})\subset V_{i}$.
If we had $cl_{U}(fV_{i})=V_{i}$ then $V_{i}$ would both open and closed in $U$, contradicting the connectivity of $U$. Hence $V_{i}-cl_{U}(fV_{i})\neq\varnothing$. Since $\mu$ is positive on open sets, we have $\mu(cl_{U}(fV_{i}))<\mu(V_{i})$. ∎
Next we present a version of Cartwright and Littlewood’s fixed point theorem, see [3].
2.24 Proposition.
Let $e$ be a fixed prime end of $U$ and $p$ a principal point of $e$. Then $f(p)=p$.
Proof.
Let $(V_{i})$ be a chain defining $e$ such that $fr_{U}(V_{i})\rightarrow p$. From Lemma 2.23 there exists $i_{0}$ such that for $i\geq i_{0}$ there exists a point $x_{i}\in fr_{U}(V_{i})$ such that $f(x_{i})$ also belongs to $fr_{U}(V_{i})$. Since $fr_{U}(V_{i})\rightarrow p$, we have $x_{i}\rightarrow p$ and $f(x_{i})\rightarrow p$ implying that $f(p)=p$. ∎
3. The structure and dynamics of the frontier of domains with fixed prime ends
From now on, $S$ will be a connected compact orientable surface and $f:S\rightarrow S$ orientation preserving $C^{1}$ diffeomorphism of $S$ such that $f^{*}\eta=\eta$, for some non degenerate $2$-form on $S$. We will still denote by $\mu$ the invariant measure induced by $\eta$. $U$ will be an $f$-invariant domain of $S$ such that $fr_{S}U$ is compact and has a finite number of connected components.
Let $p$ be a fixed point of $f$.
We say that $p$ is non degenerate if $1$ is not an eigenvalue of $df_{p}$.
A non degenerate fixed point $p$ is of saddle type or just a saddle if the eigenvalues of $df_{p}$ are real numbers $\lambda$ and $\lambda^{-1}$.
The stable and unstable invariant manifolds of $p$ are defined as:
$W_{p}^{s}=\{x\in S\thinspace|\thinspace\lim_{n\rightarrow\infty}f^{n}x=p\}$ and
$W_{p}^{u}=\{x\in S\thinspace|\thinspace\lim_{n\rightarrow-\infty}f^{n}x=p\}$, respectively.
$W_{p}^{s}$ and $W_{p}^{u}$ are injectively immersed connected curves. The components of $W_{p}^{s}-\{p\}$ and $W_{p}^{u}-\{p\}$, with respect to the topology induced by a parameterization, are called branches. By a connection we mean a branch which is contained in the intersection of two invariant manifolds (possibly of two different fixed points).
Let $p$ be a saddle of $f$ and $\mathbb{V}$ be the collection of all open neighborhoods $V$ of $p$, where there exist continuous coordinates with $p$ at the origin and in which $f(x,y)=(\lambda x,\lambda^{-1}y)$, where $\lambda,\lambda^{-1}\in\mathbb{R}$.
The components of $W_{p}^{s}\cap V$ and $W_{p}^{u}\cap V$ which contain $p$ are called the local invariant manifolds of $p$ defined by $V$. The local branches of $p$ defined by $V$ are the components of the complement of $p$ in the local invariant manifolds defined by $V$.
The set $\widetilde{V}=\{(x,y)\in V\mid x\neq 0\text{ and }y\neq 0\}$ and its components are open in $S$. There are four components of $\widetilde{V}$ that contain $p$ in their closures in $S$. We will call these sets the four sectors of p defined by V and will use $\Sigma_{i}(V)$, $1\leq i\leq 4$, to denote them.
$\Sigma_{1}(V)$ will be the component where points have coordinates $x>0$ and $y>0$, $\Sigma_{2}(V)$ the component where $x<0$ and $y>0$, $\Sigma_{3}(V)$ the component where $x<0$ and $y<0$ and $\Sigma_{4}(V)$ the component where $x>0$ and $y<0$.
If $V_{1},V_{2}\in\mathbb{V}$ then $\Sigma_{i}(V_{1})\cap\Sigma_{j}(V_{2})=\varnothing$ if $i\neq j$, or $\Sigma_{i}(V_{1})$ $\sim_{p}$ $\Sigma_{i}(V_{2})$ define the same germ at $p$.
A sector of $p$, denoted by $\Sigma_{i}(p)$ or just $\Sigma_{i}$, is the germ at $p$ of the sectors of p defined by V, $\Sigma_{i}(V)$, where $V\in\mathbb{V}$. $p$ has four sectors, $\Sigma_{i}(p)$, $1\leq i\leq 4$. Sometimes we will simplify writing by referring to the sectors through their representatives.
A set $\mathcal{A}$ contains a sector $\Sigma_{i}(p)$, if $\mathcal{A}$ contains a sector $\Sigma_{i}(V)$ that represents $\Sigma_{i}(p)$, for some $V\in\mathbb{V}$.
Now we would like to say what it means for a set $\mathcal{B}$ to accumulate on $p$ with points coming through a sector $\Sigma_{i}(p)$. A set $\mathcal{B}$ accumulates on a sector $\Sigma_{i}(p)$, if $\big{(}cl_{S}\mathcal{B}\big{)}\cap\big{(}\Sigma_{i}(V)\big{)}\neq\varnothing$ for every $V\in\mathbb{V}$.
We say that a branch $L$ and a sector $\Sigma$ are adjacent if a local branch of $L$ is contained in the closure of $\Sigma$ in $S$. Two branches are adjacent if they are adjacent to a single sector. Two sectors are adjacent if a local branch is contained within the closure of each.
We are going to use the notation ${o}(x)=o(x,f)$ for the orbit of $x$ under $f$$\{f^{n}{(x)}\in S\mid n\in\mathbb{Z}\}$.
For a branch or invariant manifold $L$ and $q_{1},q_{2}\in L$ we will denote the closed arc of points of $L$ between $q_{1}$ and $q_{2}$ by $L[q_{1},q_{2}]$. The open and half-open arcs will be denoted by $L(q_{1},q_{2})$, $L[q_{1},q_{2})$ and $L(q_{1},q_{2}]$, respectively.
We say that a fixed point $p$ is elliptic if $df_{p}$ is a rotation by an angle $\theta\neq 0$. If $p$ is a point of period $\tau$ we use $f^{\tau}$ to define these concepts for $p$.
In this section we will prove the following result.
3.1 Theorem.
Let $S$ be a compact connected orientable surface and $f:S\rightarrow S$ an area preserving orientation preserving $C^{1}$ diffeomorphism of $S$. Assume that $U$ is an invariant domain of $S$ such that $fr_{S}{U}$ has a finite number of connected components.
Let $b$ be a regular ideal boundary point of $U$ such that $f_{*}(b)=b$ and $\hat{f}:C(b)\rightarrow C(b)$ the homeomorphism on the corresponding circle of prime ends. Suppose that all fixed points of $f$ in $Z(b)$ are non degenerate.
Assume that there exists a fixed prime end $e\in C(b)$. Then we know the following.
  1. (1)
    If $p$ is the principal point of $e$ then $p$ is also a fixed point of $Z(b)$ and $p$ is a saddle.
  2. (2)
    The mapping $\hat{f}:C(b)\rightarrow C(b)$ is orientation preserving and has a finite number of fixed points.
  3. (3)
    There exists a finite singular covering $\phi:C(b)\rightarrow Z(b)$, which is a semiconjugacy between the mapping of prime ends on $C(b)$ and the restriction of $f$ to $Z(b)$.
  4. (4)
    The impression $Z(b)$ is the connected union of a finite number of saddle connections and the corresponding saddles.
3.1. The principal point of an accessible fixed prime end is a saddle
In this subsection we prove item $(1)$ of Theorem 3.1.
If $p$ is a principal point of a fixed prime end $e\in C(b)$, then by Proposition 2.24 we know that $f(p)=p$. Our hypothesis that fixed points of $f$ contained in $Z(b)$ are non degenerate implies that they are isolated fixed points. Since the principal set of $e$ is connected and all principal points are fixed, we have that $X(e)$ consists of a one point set $\{p\}$ and $e$ is accessible.
We want to prove that $p$ is a saddle. We will assume that $p$ is elliptic and obtain a contradiction.
We will write coordinates in the plane as a complex numbers. Let $V$ be a neighborhood of $p$ in $S$ where there exist coordinates with $p$ at the origin and in which $f^{\prime}(0)=e^{i\alpha}$ with $0<\alpha<2\pi$.
The idea resembles the role of the minimal invariant circles of a Moser stable fixed point in $fr_{S}U$. We consider a chain $(V_{i})$ that represents $e$ and such that $\beta_{i}=fr_{U}V_{i}$ are arcs of circles with center at 00.
If the arcs are very small, then the dynamics of $f$ and of $f^{\prime}(0)$ are close enough, so that a finite number of arcs $f^{k}(\beta_{i})$ rotate around $p$ and satisfy $f^{k}(\beta_{i})\cap f^{k+1}(\beta_{i})\neq\varnothing$ for $10\leq k\leq n-$ 1 and $f^{n}(\beta_{i})\cap\beta_{i}\neq\varnothing$. In this way they close a curve contained in $U$ with $p$ inside it.
Then it is easy to get a contradiction.
3.1.1. Estimating the distance from iterates of $z$ under $f$ and $f^{\prime}(0)$
Let $B_{\delta}^{\circ}$ be the open ball of radius $\delta$ centered at that origin with $\delta$ small enough so that $B_{\delta}^{\circ}\subset V$ and $f(B_{\delta}^{\circ})\subset V$. For points $z$ close to 00 we will need to estimate the distance between iterates of $z$ under $f$ and $f^{\prime}(0)$.
3.2 Lemma.
Let $w$ and $w^{\prime}\in\mathbb{C}$ and $\epsilon<1$ be such that $\left|w^{\prime}-w\right|<\epsilon\left|w\right|$. If $w=re^{i\eta}$ and $r^{\prime}=\left|w^{\prime}\right|$ then there exists a unique real number $\eta^{\prime}$ such that
  1. (1)
    $w^{\prime}=r^{\prime}e^{i\eta^{\prime}}$.
  2. (2)
    $\left|r^{\prime}-r\right|<\epsilon\hskip 1.42262ptr$.
  3. (3)
    $\left|\eta^{\prime}-\eta\right|<\frac{\pi}{2}\hskip 1.42262pt\epsilon$.
Proof.
Let $B_{r}$ be the closed ball of radius $\epsilon r$ with center at $w$. Then $w^{\prime}$ belongs to the interior of $B_{r}$ and $0\notin B_{r}$.
Let $\nu$ be the angle between the ray starting at 00 through $w$ and a ray starting at 00 and tangent to $B_{r}$. Since $0\notin B_{r}$ we have $\nu<\frac{\pi}{2}$. Then there exists a unique real number $\eta^{\prime}$ such that $w^{\prime}=r^{\prime}e^{i\eta^{\prime}}$ and $\left|\eta^{\prime}-\eta\right|<\nu$.
For $0<\nu<\frac{\pi}{2}$ we have $\frac{2}{\pi}\nu<\sin\nu$. Since $\sin\nu=\epsilon$, we have $\left|\eta^{\prime}-\eta\right|<\frac{\pi}{2}\epsilon$, which proves $(1)$ and $(3)$.
$(2)$ follows from the triangle inequality. ∎
For $\epsilon\in(0,1)$, let $\delta>0$ be such that if $\left|z\right|<\delta$ then $\left|f(z)-f^{\prime}(0)z\right|<\epsilon\left|z\right|$.
3.3 Lemma.
Let $z=re^{i\theta}$ with $r<\delta$ and $r^{\prime}=\left|f(z)\right|$.
Then there exists a unique real number $\theta^{\prime}$ such that
  1. (1)
    $f(z)=r^{\prime}e^{i\theta^{\prime}}$.
  2. (2)
    $\left|r^{\prime}-r\right|<\epsilon r$.
  3. (3)
    $\left|\theta^{\prime}-(\theta+\alpha)\right|<\frac{\pi}{2}\epsilon$.
Proof.
Let $w^{\prime}=f(z)$ and $w=f^{\prime}(0)z$. Then $w=re^{i(\theta+\alpha)}$ and $\left|w^{\prime}-w\right|<\epsilon\left|w\right|$.
Let $\eta=\theta+\alpha$. Lemma 3.2 provides a unique number $\theta^{\prime}=\eta^{\prime}$ for which $(1)$ $(2)$ and $(3)$ are satisfied. ∎
By Lemma 3.3, the mapping $F:\mathbb{R}\times(0,\delta)\rightarrow\mathbb{R}\times(0,+\infty)$ defined by $F(\theta,r)=(\theta^{\prime},r^{\prime})$ is a lifting of $f:B_{\delta}-\{0\}\rightarrow\mathbb{C}$ with the following properties:
  • $\left|r^{\prime}-r\right|<\epsilon r$.
  • $\left|\theta^{\prime}-(\theta+\alpha)\right|<\frac{\pi}{2}\epsilon$
for every $(\theta,r)\in\mathbb{R}\times(0,\delta)$.
We are going to use $(\theta_{j},r_{j})=F^{j}(\theta_{0},r_{0})$ to denote iterates of a point $(\theta_{0},r_{0})$.
3.1.2. The construction
Now we start the construction.
Let $B\subset V$ be a closed ball centered at 00 such that $Z(b)$ is not contained in $B$.
Let $n$ such that $n\alpha>2\pi$ and $n(2\pi-\alpha)>2\pi$.
Let $\epsilon\in(0,1)$ such that
(2) $n\frac{\pi}{2}\epsilon<n\alpha-2\pi,$
(3) $n\frac{\pi}{2}\epsilon<n(2\pi-\alpha)-2\pi,$
(4) $\alpha+\frac{\pi}{2}\epsilon<2\pi\text{ and }\alpha-\frac{\pi}{2}\epsilon>0.$
Let $\delta>0$ be such that
  1. (1)
    $B_{\delta}\subset B^{\circ}$.
  2. (2)
    $f(B_{\delta}^{\circ})\subset V$.
  3. (3)
    If $\left|z\right|<\delta$ then $\left|f(z)-f^{\prime}(0)z\right|<\epsilon\left|z\right|$.
Suppose that the coordinate $r_{0}$ of the initial point is a positive number that satisfies $r_{0}(1+\epsilon)^{n}<\delta$.
Then from item $(2)$ of Lemma 3.3 we have $r_{j}<r_{0}(1+\epsilon)^{j}<\delta$ for $1\leq j\leq n$. Therefore, if
(5) $r_{0}<\frac{\delta}{(1+\epsilon)^{n}}$
then the iterates $(\theta_{j},r_{j})=F^{j}(\theta_{0},r_{0})$ are well defined for $1\leq j\leq n$.
From item $(3)$ of Lemma 3.3 we have that
[…]
From Lemma 2.20, we know that there exists $\delta_{1}>0$ such that for any sequence $(\rho_{i})$ contained in $(0,\delta_{1})$ such that $lim_{i\rightarrow\infty}\rho_{i}=0$, there exists a chain $(V_{i})$ representing $e$ such that $fr_{U}V_{i}\rightarrow p$ and the sets $\beta_{i}:=fr_{U}V_{i}$ are arcs of circles of radius $\rho_{i}$ and center at 00.
There exists $i_{1}$ such that $\beta_{i}\subset C_{\rho_{i}}$ with $\rho_{i}<\frac{\delta}{(1+\epsilon)^{n}}$ for $i\geq i_{1}$.
If we apply Lemma 2.23 to $f$ and $f^{n}$ at the same time, we conclude that there exists $i_{2}$ such that $f(\beta_{i})\cap\beta_{i}\neq\varnothing$ and $f^{n}(\beta_{i})\cap\beta_{i}\neq\varnothing$ for $i\geq i_{2}$.
Therefore, if we choose $\beta$ to be one of the arcs $\beta_{i}$ with $i\geq\max\{i_{1},i_{2}\}$, then $\beta$ satisfies the following.
  1. (1)
    If the radius of $\beta$ is $r_{0}$, then $r_{0}$ that satisfies equation 5, and the first $n$ iterates of $\beta$ by $F$ are well defined.
  2. (2)
    $f^{l}(\beta)\cap f^{l+1}(\beta)\neq\varnothing$ for $0\leq l-1\leq n$ and $f^{n}(\beta)\cap\beta\neq\varnothing$.
3.1.3. The iterates of $\beta$ turn around $p$ and close a curve in $U$
We are going to show that these arcs turn around $p$ at least once to close a curve with $p$ inside.
Let $\hat{\beta}=(a,b)\times\{r_{0}\}$ be a lifting of $\beta$ with $0\leq a<2\pi$. We have $b-a<2\pi$.
3.4 Lemma.
$F\hat{\beta}\cap\hat{\beta}\neq\varnothing$ or $F\hat{\beta}\cap(\hat{\beta}+(2\pi,0))\neq\varnothing$.
Proof.
Since $f\beta\cap\beta\neq\varnothing$ we have that $F\hat{\beta}\cap(\hat{\beta}+(2k\pi,0))\neq\varnothing$ for some $k\in\mathbb{Z}$.
We are going to show that $k=0$ or $1$.
Let $(\theta_{1},r_{1})\in F\hat{\beta}\cap(\hat{\beta}+(2k\pi,0))$. Then $(\theta_{1},r_{1})=F(\theta_{0},r_{0})$ and $\theta_{1}=\overline{\theta}_{0}+2k\pi$, where both $\theta_{0}$ and $\overline{\theta}_{0}$ belong to $(a,b)$.
By equation (6) we have $\left|\theta_{1}-(\theta_{0}+\alpha)\right|<\frac{\pi}{2}\epsilon$. Therefore
$\left|\overline{\theta}_{0}+2k\pi-(\theta_{0}+\alpha)\right|<\frac{\pi}{2}\epsilon,\hskip 14.22636pt\left|2k\pi-\alpha\right|<\left|\theta_{0}-\overline{\theta}_{0}\right|+\frac{\pi}{2}\epsilon<2\pi+\frac{\pi}{2}\epsilon$
and
$-2\pi+\alpha-\frac{\pi}{2}\epsilon<2k\pi<2\pi+\alpha+\frac{\pi}{2}\epsilon.$
By equations (4) we have $-2\pi<2k\pi<4\pi$ and $k=0$ or $1$. ∎
3.5 Lemma.
$F^{n}\hat{\beta}\cap(\hat{\beta}+(2k\pi,0))\neq\varnothing$ for some $k\geq 1$.
Proof.
Since $f^{n}\beta\cap\beta\neq\varnothing$ we have that $F^{n}\hat{\beta}\cap(\hat{\beta}+(2k\pi,0))\neq\varnothing$ for some $k\in\mathbb{Z}$.
Let $(\theta_{n},r_{n})\in F^{n}\hat{\beta}\cap(\hat{\beta}+(2k\pi,0))$. Then $(\theta_{n},r_{n})=F^{n}(\theta_{0},r_{0})$ and $\theta_{n}=\overline{\theta}_{0}+2k\pi$ where both $\theta_{0}$ and $\overline{\theta}_{0}$ belong to $(a,b)$.
By equation 6, we have $\left|\theta_{n}-(\theta_{0}+n\alpha)\right|<n\frac{\pi}{2}\epsilon$. Therefore
$\left|\overline{\theta}_{0}+2k\pi-(\theta_{0}+n\alpha)\right|<n\frac{\pi}{2}\epsilon,\hskip 14.22636pt2k\pi+\overline{\theta}_{0}-\theta_{0}-n\alpha>-n\frac{\pi}{2}\epsilon$
and
$2k\pi>\theta_{0}-\overline{\theta}_{0}+n\alpha-n\frac{\pi}{2}\epsilon>-2\pi+n\alpha-n\frac{\pi}{2}\epsilon>0,$
by equation (2).
Hence $k\geq 1$. ∎
3.1.4. The proof when $F(\hat{\beta})$ intersects $\hat{\beta}$
As we have seen in Lemma 3.4, we have two possibilities: $F\hat{\beta}\cap\hat{\beta}\neq\varnothing$ or $F\hat{\beta}\cap(\hat{\beta}+(2\pi,0))\neq\varnothing$.
We will finish proving item $(1)$ of Theorem 3.1 in the case where $F\hat{\beta}\cap\hat{\beta}\neq\varnothing$. We construct $\xi$ in the following way.
We have $F^{j+1}\hat{\beta}\cap F^{j}\hat{\beta}\neq\varnothing$ for $10\leq j\leq n-$ 1 and by Lemma 3.5 $F^{n}\hat{\beta}\cap(\hat{\beta}+(2k\pi,0))\neq\varnothing$ for some $k\geq 1$.
This implies that the union $\hat{\beta}\cup F\hat{\beta}\cup\ldots\cup F^{n}\hat{\beta}\cup(\hat{\beta}+(2k\pi,0))$ is path connected and therefore this union contains a path $\hat{\xi}$ connecting any point $(\theta_{0},r_{0})\in\hat{\beta}$ to $(\theta_{0}+2k\pi,r_{0})$ for some $k\geq 1$.
Let $P:\mathbb{R}\times(0,\infty)\rightarrow\mathbb{R}^{2}\setminus\{(0,0)\}$ be the covering map and $\xi=P(\hat{\xi})$.
3.6 Lemma.
$\xi$ is a closed path of $S$ that satisfies the following.
  1. (1)
    $\xi\subset U$.
  2. (2)
    $\xi\subset B_{\delta}^{\circ}-\{0\}$.
  3. (3)
    If $B^{\prime}$ is the component of $S\setminus\xi$ which contains $p=0$ then
    1. (a)
      $B^{\prime}\subset B^{\circ}$.
    2. (b)
      $fr_{S}B^{\prime}\subset\xi$.
Proof.
Since $\beta\subset U$ and $\xi\subset\cup_{j=0}^{n}f^{j}\beta$ we have that $\xi\subset U$, which proves item $(1)$.
Item $(2)$ follows immediately from the construction of $\xi$. Observe that $\xi$ is a non trivial element of $\pi_{1}(B_{\delta}^{\circ}-\{0\})$.
Let $B^{\prime}$ be the component of $S\setminus\xi$ which contains $p$.
If $\delta_{1}<\inf\{r>0\mid(\theta,r)\in\hat{\xi}\text{ for some }\theta\in[\theta_{0},\theta_{0}+2k\pi]\}$ then $P^{-1}(\beta^{\prime})$ contains ×$\mathbb{R}\times(0,\delta_{1})$.
Similarly, if $\delta>\delta_{2}>\sup\{r>0\mid(\theta,r)\in\hat{\xi}\text{ for some }\theta\in[\theta_{0},\theta_{0}+2k\pi]\}$ then $P^{-1}(\beta^{\prime})$ is contained ×$\mathbb{R}\times(0,\delta_{2})$.
This implies that $B_{\delta_{1}}^{\circ}\subset B^{\prime}\subset B_{\delta_{2}}^{\circ}$. Therefore $B^{\prime}\subset B_{\delta}^{\circ}\subset B^{\circ}$, which proves item $(3a)$. Item $(3b)$ is obvious. ∎
The contradiction is obtained in the following way.
Since $X(e)=\{p\}\subset Z(b)$ we have that $Z(b)\cap B^{\prime}\neq\varnothing$. On the other hand, since $B^{\prime}\subset B$ and $Z(b)$ is not contained in $B$, we have that $Z(b)\cap(S-B^{\prime})\neq\varnothing$.
Being connected, $Z(b)\cap fr_{S}{B^{\prime}}\neq\varnothing$ and therefore $fr_{S}{B^{\prime}}\cap fr_{S}U\neq\varnothing$.
On the other hand, by Lemma 3.6, we have $fr_{S}B^{\prime}\subset\xi\subset U$.
3.1.5. The proof when $F(\hat{\beta})$ intersects $\hat{\beta}$$+(2\pi,0)$
When $F\hat{\beta}\cap(\hat{\beta}+(2\pi,0))\neq\varnothing$ we work with a different lifting of $f$, $H(\theta,r)=F(\theta,r)-(2\pi,0)$.
3.7 Lemma.
$H$ satisfies $H\hat{\beta}\cap\hat{\beta}\neq\varnothing$ and $H^{n}\hat{\beta}\cap(\hat{\beta}+(2k\pi,0))\neq\varnothing$ for some $k\leq-1$.
Proof.
$H\hat{\beta}\cap\hat{\beta}\neq\varnothing$ follows immediately from the definition of $H$ and the fact that $F\hat{\beta}\cap(\hat{\beta}+(2\pi,0))\neq\varnothing$.
$f^{n}\beta\cap\beta\neq\varnothing$ implies that $H^{n}\hat{\beta}\cap(\hat{\beta}+(2k\pi,0))\neq\varnothing$ for some $k\in\mathbb{Z}$.
Let $q\in H^{n}\hat{\beta}\cap(\hat{\beta}+(2k\pi,0))$. Since $H^{n}(\theta_{0},r_{0})=F^{n}(\theta_{0},r_{0})-(2n\pi,0)$ we have that
$q=F^{n}(\theta_{0},r_{0})-(2n\pi,0)=(\theta_{n},r_{n})-(2n\pi,0)\text{ and }q=(\overline{\theta}_{0}+2k\pi,r_{0}),$
where both $\theta_{0}$ and $\overline{\theta}_{0}$ belong to $(a,b)$.
By equation (6), we have $\left|\theta_{n}-(\theta_{0}+n\alpha)\right|<n\frac{\pi}{2}\epsilon$. Therefore
$\left|\overline{\theta}_{0}+2k\pi+2n\pi-\theta_{0}-n\alpha\right|<n\frac{\pi}{2}\epsilon,\hskip 14.22636pt2k\pi+\overline{\theta}_{0}-\theta_{0}+n(2\pi-\alpha)<n\frac{\pi}{2}\epsilon,$
$02k\pi<\theta_{0}-\overline{\theta}_{0}-n(2\pi-\alpha)+n\frac{\pi}{2}\epsilon<2\pi-n(2\pi-\alpha)+n\frac{\pi}{2}\epsilon<$ 0
by equation (3). This implies that $k\leq-1.$
Now we have $H\hat{\beta}\cap\hat{\beta}\neq\varnothing$ and $H^{n}\hat{\beta}\cap(\hat{\beta}+(2k\pi,0))\neq\varnothing$ for some $k\leq-1$. The construction of $\xi$ is done as in the previous subsubsection and we arrive at a contradiction in the same way.
This proves item $(1)$ of Theorem 3.1.
3.2. The covering $\phi:C(b)\rightarrow Z(b)$
In this subsection we will prove items $(2)$, $(3)$ and $(4)$ Theorem 3.1.
Recall our assumptions. $S$ is a compact connected orientable surface and $f:S\rightarrow S$ is an area preserving orientation preserving $C^{1}$ diffeomorphism of $S$. $U$ is an invariant domain of $S$ such that $fr_{S}{U}$ has a finite number of connected components. $b$ is a regular ideal boundary point of $U$ such that $f_{*}(b)=b$, $\hat{f}:C(b)\rightarrow C(b)$ is the homeomorphism on the related circle of prime ends, and all fixed points of $f$ in $Z(b)$ are non degenerate.
We assume that there exists a fixed prime end $e\in C(b)$.
In the previous subsection, we saw that if $e\in C(b)$ is fixed prime end then $e$ is accessible and its principal point $p$ is an isolated fixed point of $Z(b)$ of saddle type.
3.2.1. The accumulation lemma
Now we will present a very useful result in conservative dynamics on surfaces. For a proof see Corollary 8.3 of [13].
3.8 Proposition.
The Accumulation Lemma. Let $S$ be a connected surface, $f:S\rightarrow S$ an area preserving homeomorphism of $S$ and $K$ a compact connected invariant set of $f$. If $L$ is a branch of $f$ and $L\cap K\neq\varnothing$ then $L\subset K$.
Next, we would like to prove a corollary of the Accumulation Lemma that will be needed later.
3.9 Corollary.
Let $f:S\rightarrow S$ be an area preserving $C^{1}$ diffeomorphism of a compact connected surface $S$ and $L$ be a branch of a saddle fixed point $p$. Then we have the following:
  1. (1)
    Let $U$ be an invariant domain such that $fr_{S}U$ has finitely many components. If $L\cap U\neq\varnothing$ then $U$ contains $L$ and the two sectors of $p$ adjacent to $L$.
  2. (2)
    If $K$ is a compact connected invariant set then either $L\subset K$ or $L$ and its adjacent sectors are contained in one component of $S-K$.
Proof.
Assume by contradiction that $L\cap U\neq\varnothing$. If $L$ were not contained in $U$, then $L$ would intersect a component of $fr_{S}U$, which is a periodic set. By the accumulation Lemma applied to a power of $f$, this component would contain $L$, a contradiction. Therefore $L\subset U$.
Let $x\in L$ and $W\subset U$ be a neighborhood in $U$ of the arc from $x$ to $f^{2}(x)$ in $L$. Then $\cup_{n\in\mathbb{Z}}f^{2n}(W)$ contains the sectors of $p$ adjacent to $L$.
This proves item $(1)$. $(2)$ follows from $(1)$. ∎
3.2.2. The local semiconjugacy around a fixed prime end
We start by taking continuous coordinates $(x,y)$ in a neighborhood $V$ of $p$ with $p$ at the origin, where $f(x,y)=(\lambda x,\lambda^{-1}y)$ with $|\lambda|>1$. We may assume that $V$ is the open ball $B_{1}^{\circ}$ of radius $1$ and center at $(0,0)$.
There exists a path $\beta:(0,1]\rightarrow U$ such that $\underset{t\rightarrow 0}{\lim}\beta(t)=p$ in $S$ and $\underset{t\rightarrow 0}{\lim}\beta(t)=e$ in $E(U)$. We may assume that $\beta(0,1)\subset V$ and $\beta(1)\notin V$.
From item Lemma 2.20, there exists $\delta\in(0,1)$ such that for any decreasing sequence $(r_{n})_{n\geq 1}$ contained in $(0,\delta)$ with $\underset{n\rightarrow\infty}{\lim}r_{n}=0$, there exists a chain $(V_{n})$ representing $e$ such that $\xi_{n}=fr_{U}V_{n}\subset C_{r_{n}}$.
Since $\underset{t\rightarrow 0}{\lim}\beta(t)=e$ in $E(U)$ we may assume that $\xi_{n}\cap\beta\neq\varnothing$ for every $n$, and therefore $\beta(0,1)\cup(\cup_{n}\xi_{n})$ is connected and contained in both $U$ and $V$. Therefore $\beta$ and all the arcs $\xi_{n}$ must be contained in one connected component of $U\cap V$.
Denote this component by $W$.
By Corollary 3.9 a branch of $p$ is either contained in $U$ or is disjoint from it.
Note that our hypotheses do not allow taking powers of $f$.
3.10 Lemma.
We have that $\lambda>1$.
Proof.
$W$ can not contain the four local branches, otherwise $Z(b)=\{p\}$, contradicting that $b$ is regular. From this we conclude that one of the local branches is not contained in $W$ and therefore this branch is disjoint from $W$ and from $U$.
Assume by contradiction that $\lambda<{-1}$. In this case all local branches have period two.
If $U$ is disjoint from, say, $L_{1}=\{(x,y)\in V\mid x>0,y=0\}$ then it is disjoint from the local invariant manifold, $\{(x,y)\in V\mid-1<x<1,y=0\}$. This local invariant manifold separates $V$ into two ”half open” balls, $B_{+}$ and $B_{-}$.
If $\beta$ is contained in $B_{+}$ then $f\circ\beta$ is contained in $B_{-}$. As $t\rightarrow 0$ both paths tend to $p$ in $S$, but $\beta$ tends to $e$ in $E(U)$ and $f\circ\beta$ tends to $\hat{f}(e)$.
But since $\beta$ is contained in $B_{+}$ and $f\circ\beta$ is contained in $B_{-}$, we have that $\hat{f}(e)\neq e$ and $\hat{f}^{2}(e)=e$, contradicting the fact that $e$ is a fixed prime end. ∎
3.11 Lemma.
For each fixed prime end $e\in C(b)$ there exists an arc of prime ends $(a_{1},a_{2})$ around $e$ in $C(b)$ such that $\phi_{e}:(a_{1},a_{2})\rightarrow Z(b)$ defined by $\phi_{e}(a)=Y(a)$ is a function that satisfies the following:
  1. (1)
    $Y(a)$ is a one point set for every $a\in(a_{1},a_{2})$.
  2. (2)
    $\phi_{e}(e)=p$ and $\phi_{e}((a_{1},e))$ and $\phi_{e}((e,a_{2}))$ are local branches of $p$.
  3. (3)
    $\big{(}\phi_{e}\circ\hat{f}\big{)}(a)=\big{(}f\circ\phi_{e}\big{)}(a)$ if $a$ and $\hat{f}(a)$ belong to $(a_{1},a_{2})$.
Proof.
Let us first consider the case in which $W$ does not intersect any of the branches of $p$. In this case $W$ must be contained in one of the four sectors of $p$ defined by $V$, say $S_{1}=\{(x,y)\in V\hskip 1.42262pt|\hskip 1.42262ptx>0,y>0\}$. The local branches adjacent to $S_{1}$ are $L_{1}$ and $L_{2}=\{(x,y)\in V\mid x=0,y>0\}$.
From Lemma 2.23, there exists $n_{0}$ such that $\xi_{n}\cap f(\xi_{n})\neq\varnothing$ for $n\geq n_{0}$.
Lemma 2.20 let us choose any norm in $\mathbb{R}^{2}$ to work with. We are going to consider the sup norm, for which a ball $B_{\rho}$ is the square with vertices at (±$(\pm\rho,\pm\rho)$.
Since $(r_{n})$ is an arbitrary sequence contained in $(0,\delta)$, given any number $\rho\in(0,r_{n_{0}})$, we may assume that the sequence $(r_{n})$ provided by Lemma 2.20 satisfies $\rho=r_{m}$ for some $m>n_{0}$.
Let $\Gamma_{\rho}=\big{(}(0,\rho]\times\{\rho\})\cup(\{\rho\}\times(0,\rho]\big{)}$. Then $\xi_{m}\subset\Gamma_{\rho}$, and since $f(\Gamma_{\rho})\cap\Gamma_{\rho}=\{(\rho,\lambda^{-1}\rho)\}$ we have that $f(\xi_{m})\cap\xi_{m}=\{(\rho,\lambda^{-1}\rho)\}$ as well.
Therefore $(\rho,\lambda^{-1}\rho)$ and $(\lambda^{-1}\rho,\rho)$ belong to $\xi_{m}$, and the arc $\Lambda_{\rho}$ from $(\rho,\lambda^{-1}\rho)$ to $(\lambda^{-1}\rho,\rho)$ inside $\Gamma_{\rho}$ is contained in $\xi_{m}\subset U$. This holds for every $\rho\in(0,r_{n_{0}})$.
The set $C:=\underset{\rho\in(0,r_{n_{0}})}{\bigcup}\Lambda_{\rho}$ is the intersection of $B^{\circ}_{r_{n_{0}}}$ with the closed cone with vertex at $(0,0)$ and boundary at the half-lines from $(0,0)$ and slopes $\lambda^{-1}$ and $\lambda$.
We have that $C\subset U$, and therefore $R:=(0,r_{n_{0}})\times(0,r_{n_{0}})\subset\underset{n\in\mathbb{Z}}{\bigcup}f^{n}C\subset U$.
Since $R\subset V$ and $R$ intersects the arcs $\xi_{m}$, we conclude that $R\subset W$.
For any $q\in\Big{(}\big{(}[0,r_{n_{0}})\times\{0\}\big{)}\cup\big{(}\{0\}\times[0,r_{n_{0}})\big{)}\Big{)}$, if $B_{\frac{1}{n}}^{\circ}(q)$ is the open ball of radius $\frac{1}{n}$ with center at $q$, then $V_{n}=B_{\frac{1}{n}}^{\circ}(q)\cap R$ defines a prime chain $(V_{n})=a$ such that $Y(a)=\{q\}$.
Let $(a_{1},a_{2})$ be the arc of prime ends whose impressions are the points of $\big{(}[0,r_{n_{0}})\times\{0\}\big{)}\cup\big{(}\{0\}\times[0,r_{n_{0}})\big{)}$. We define $\phi_{e}(a)=Y(a)$ for $a\in(a_{1},a_{2})$.
Items $(1)$ and $(2)$ are easy to check.
Item $(3)$ follows from the general fact that for any $e^{\prime}\in E(U)$ we have the equality of sets $f\big{(}Y(e^{\prime})\big{)}=Y\big{(}\hat{f}(e^{\prime})\big{)}$.
This completes the construction of the local semiconjugacy when $W$ does not intersect any branch of $p$.
We will need item $(1)$ of Corollary 3.9, which says that if $L$ is a branch of $p$ and $L\cap U\neq\varnothing$ then $U$ contains $L$ and the two sectors of $p$ adjacent to $L$.
Now we consider the case where $W$ intersects a branch of $p$, say the branch that contains the local branch $L_{2}$, and $W$ does not intersect the branches that contain $L_{1}$ and $L_{3}=\{(x,y)\in V\hskip 1.42262pt|\hskip 1.42262ptx<0,y=0\}$.
From Corollary 3.9, we have that $W=S_{1}\cup L_{2}\cup S_{2}$, where $S_{2}=\{(x,y)\in V\hskip 1.42262pt|\hskip 1.42262ptx<0,y>0\}$. It follows that, any $q\in\{(x,y)\in V\hskip 1.42262pt|\hskip 1.42262pt-1<x<1,y=0\}$ is the impression of a prime end defined by chains made of ”half” open balls centered at $q$ and contained in $W=S_{1}\cup L_{2}\cup S_{2}$. The semiconjugacy is constructed as in the previous case.
Now it should be clear that there are four possibilities for $W$: $W$ consists one sector; $W$ consists of a local branch and its two adjacent sectors; $W$ consists of three sectors and the two local branches between them; $W$ consists of the complement in $V$ of $p$ and and one local branch.
As we said before, $W$ can not contain the four local branches.
We described the construction of the local semiconjugacy in the first two cases. In the other two, the construction is analogous to the second case. ∎
3.2.3. The construction of the semi conjugacy
The existence of the local semiconjugacy $\phi_{e}:(a_{1},a_{2})\rightarrow Z(b)$ around every fixed prime end $e$ of $C(b)$ implies that they are isolated and exist in a finite number.
Since $\phi_{e}(e)=p$ and each branch of $p$ is invariant, the existence of the local semiconjugacy also implies that $(a_{1},e)$ and $(e,a_{2})$ are mapped by $\hat{f}$ onto arcs that satisfy $\hat{f}(a_{1},e)\cap(a_{1},e)\neq\varnothing$ and $\hat{f}(e,a_{2})\cap(e,a_{2})\neq\varnothing$.
From this we conclude that $\hat{f}:C(b)\rightarrow C(b)$ is orientation preserving and has a finite number of fixed points. This proves item $(2)$ of Theorem 3.1.
Let $[e_{1},e_{2}]$ be an arc of prime ends, where the end points are fixed, and if $e\in(e_{1},e_{2})$ then $\underset{n\rightarrow\infty}{lim}\hat{f}^{n}(e)=e_{2}$ and $\underset{n\rightarrow-\infty}{lim}\hat{f}^{n}(e)=e_{1}$. For $i=1,2$ let $(a_{i1},a_{i2})$ be the arc around $e_{i}$ where the local semiconjugacy $\phi_{e_{i}}:(a_{i1},a_{i2})\rightarrow Z(b)$, $\phi_{e_{i}}(a)=Y(a)$, is defined.
Let $a\in(e_{1},e_{2})$. There exist $n_{1}$ and $n_{2}$ such that $f^{n_{i}}(a)\in(a_{i1},a_{i2})$. We have that $f^{n}{\big{(}Y(e^{\prime})\big{)}}=Y\big{(}(\hat{f})^{n}(e^{\prime})\big{)}$ for any $e^{\prime}\in E(U)$ and $n\in\mathbb{Z}$. Since $Y\big{(}\hat{f}^{n_{i}}(a)\big{)}$ is a one point set, the same happens to $Y(a)$.
Therefore $Y(a)$ is a one point set for every $a\in C(b)$ and $\phi:C(b)\rightarrow Z(b)$, $\phi(a)=Y(a)$, is well defined.
Obviously $f\circ\phi=\phi\circ\hat{f}$ on $C(b)$.
Each arc $(a_{i1},a_{i2})$ is mapped into the union of $p_{i}=Y(e_{i})$ and two of its local branches. Since $f^{n_{i}}(a)\in(a_{i1},a_{i2})$ for $i=1,2$, we have that $Y(a)$ belongs to a branch of each $p_{i}$.
From this we conclude that the arc $(e_{1},e_{2})$ is mapped onto a connection from $p_{1}$ to $p_{2}$.
Now we show that $\phi$ is continuous. Let $a\in C(b)$.
Consider a neighborhood $A\cap Z(b)$ of $\phi(a)=Y(a)$ in $Z(b)$, where $A$ is an open subset of $S$. If $a$ is represented by a prime chain $(V_{n})$ then $Y(a)=\cap_{n}cl_{S}{V_{n}}$, and therefore $cl_{S}{V_{n}}\subset A$ for some $n$. We have that $V^{\prime}_{n}$ is a neighborhood of $a$ in $E(U)$. If $e\in C(b)\cap V^{\prime}_{n}$, then $e=(W_{k})$ with $W_{k}\subset V_{n}$ for some $k$.
It follows that $Y(e)=\cap_{k}cl_{S}{W_{k}}\in A\cap Z(b)$, and $\phi$ is continuous at $a$.
It remains to show that $\phi:C(b)\rightarrow Z(b)$ is surjective. But this follows from Proposition 2.19 which says that $Z(b)=\cup_{e\in C(b)}Y(e)$.
This proves items $(3)$ and $(4)$ of Theorem 3.1 and its proof is complete.
Finally, we would like to present a result of Mather under weaker hypotheses.
3.12 Proposition.
Let $S$ be a connected surface, $f:S\rightarrow S$ an area preserving orientation preserving $C^{1}$ diffeomorphism and $K$ be a compact connected invariant set of $f$ such that every fixed point of $f$ in $K$ is non degenerate.
Let $p$ be a saddle fixed point of $f$ and assume that no branch of $p$ is a connection.
If $p\in K$ then $W_{p}^{u}\cup W_{p}^{s}\subset K$.
Proof.
Assume by contradiction that a branch $L$ of $p$ is not contained in $K$.
By item $(2)$ of Corollary 3.9, $L$ and its adjacent sectors are contained in one component $U$ of $S-K$.
We know that $fr_{S}U$ is compact and has a finite number of connected components. Since $L$ is invariant, $U$ is also invariant. If we parameterize the local branch of $L$ by $\alpha:(0,1]\rightarrow L$ so that $\underset{t\rightarrow 0^{+}}{\lim}\alpha(t)=p$ then, as $t\rightarrow 0^{+}$, we have that $\alpha(t)\rightarrow e$ for some $e\in E(U)$ and $\alpha(t)\rightarrow b$ for some $b\in B(U)$.
Since $Z(b)\subset fr_{S}U\subset K$, all fixed points of $f$ in $Z(b)$ are non degenerate. We have that $e$ is a fixed point of $\hat{f}:C(b)\rightarrow C(b)$ and that $X(e)=\{p\}\subset Z(b)$.
By Theorem 3.1 there exists a finite to one semiconjugacy $\phi:C(b)\rightarrow Z(b)$ and $Z(b)$ is a finite union of connections that contains at least one branch of $p$.
This contradicts the fact that no branch of $p$ is a connection. ∎
3.3. Some examples and simple consequences of Theorem 3.1
3.13 Corollary.
Let $S$ be a compact connected orientable surface and $f:S\rightarrow S$ an area preserving orientation preserving $C^{1}$ diffeomorphism of $S$. Assume that $U$ is a periodic domain of $S$ such that $fr_{S}{U}$ has a finite number of connected components. Suppose that all periodic points of $f$ in $fr_{S}U$ are non degenerate.
Let $E(U)$ be the prime ends compactification of $U$ and assume that all induced maps on circles of prime ends have rational rotation number.
Then $fr_{S}U$ is the union of finitely many saddle connections and the corresponding saddles.
3.14 Corollary.
Under the same hypotheses of Corollary 3.13, if $U$ is homeomorphic to an open disk then $fr_{S}U$ is the connected union of finitely many saddle connections and the corresponding saddles.
3.15 Example.
Let $H:\mathbb{R}^{2}\rightarrow\mathbb{R}$ given by $H(x,y)=\sin{(\pi x)\sin{(\pi y)}}$.
We have that $H(x+2k,y+2l)=H(x,y)$ for every $(k,l)\in\mathbb{Z}^{2}$ and $H$ is periodic in both variables with period $2$.
Consider the Hamiltonian flow generated by $H$ and let $f:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ be its time one map. We know that $f$ preserves the two dimensional Lebesgue measure.
If we look at the Hamiltonian equations of the flow and the level curves of $H$, then we conclude the following.
  1. (1)
    $f(x+2k,y+2l)=f(x,y)$ for every $(k,l)\in\mathbb{Z}^{2}$.
  2. (2)
    All fixed points are saddles or elliptic.
  3. (3)
    A fixed point $p$ is a saddle if and only if $p\in\mathbb{Z}^{2}$.
  4. (4)
    A fixed point is elliptic if and only if $(x,y)=(k+\frac{1}{2},l+\frac{1}{2})$ for $(k,l)\in\mathbb{Z}^{2}$.
  5. (5)
    The segments $(k,k+1)\times\{l\}$ and $\{k\}\times(l,l+1)$, where $(k,l)\in\mathbb{Z}^{2}$ are all saddle connections.
With these properties in mind we will construct a couple of examples on the torus $\mathbb{T}^{2}$.
We will think of $\mathbb{T}^{2}$ as the square with vertices at (±2,±2($\pm$ 2$\pm$ 2( ± 2 , ± 2), with opposite sides identified. We will still denote by $f$ as the induced map on $\mathbb{T}^{2}$.
In the first example we consider $K=\big{(}[-1,1]\times\{0\}\big{)}\cup\big{(}\{0\}\times[-1,1]\big{)}$ and $U=\mathbb{T}^{2}\setminus K$. We have that $U$ is a surface of genus one that has only one ideal boundary point, $b(U)=\{b\}$, and $B(U)$ is homeomorphic to a torus.
We have that $K$ and $U$ are invariant under $f$. $K=fr_{S}U=Z(b)$ is made of $5$ fixed points of saddle type, $(0,0)$, (±$(\pm 1,0)$ and $(0,\pm 1)$, and $4$ connections, $(-1,0)\times\{0\}$, $(0,1)\times\{0\}$, $\{0\}\times(-1,0)$ and $\{0\}\times(0,1)$.
On the other hand, $E(U)$ is a surface of genus one and one boundary component $C(b)$. The restriction of $\hat{f}$ to $C(b)$ has $8$ fixed points. If we group these fixed points into two groups of 4 points each, so that in a circular order the points of each group alternate, then the semiconjugacy $\phi:C(b)\rightarrow Z(b)$ takes one group into $(0,0)$, and the other 4 fixed points into each one of the remaining saddles of $K$, (±$(\pm 1,0)$ and $(0,\pm 1)$.
Each saddle connection is the image of two arcs between fixed points of $C(b)$.
Now we describe the second example.
Let $C$ be the circle $[-2,2]\times\{0\}$, $I_{1}=\{0\}\times[0,1]$, $I_{2}=\{0\}\times[-1,0]$, $K=C\cup I_{1}\cup I_{2}$ and $U=\mathbb{T}^{2}\setminus K$.
We have that $K$ is compact and connected, $K=fr_{S}U$, $U$ has two ideal boundary points, $b(U)=\{b_{1},b_{2}\}$ and $B(U)$ is homeomorphic to a sphere. If we approach $C\cup I_{1}$ using points of $U$ above $C$ and call this ideal boundary point $b_{1}$ then $Z(b_{1})=C\cup I_{1}$. Approaching from below we call the ideal boundary point $b_{2}$ and $Z(b_{2})=C\cup I_{2}$. Both $Z(b_{1})$ and $Z(b_{2})$ have $5$ saddles and $5$ connections.
We have that $E(U)$ is a compact surface of genus zero and two boundary components, $C(b_{1})$ and $C(b_{2})$. Each circle of prime ends has $6$ fixed points of $\hat{f}$.
Each semi conjugacy $\phi_{i}:C(b_{i})\rightarrow Z(b_{i})$ maps two fixed points of $C(b_{i})$ into $(0,0)$ and the other fixed points of $C(b_{i})$ are mapped into one fixed point of $Z(b_{i})$ each. $I_{i}$ contains the only connection of $Z(b_{i})$ which is the image of two arcs between fixed points of $C(b_{i})$.
4. The accumulation of branches and homoclinic points
In this section we prove some results about the accumulation of invariant manifolds and the existence of homoclinic points in low genus.
4.1. The accumulation of branches
4.1 Proposition.
Let $S$ be a connected compact orientable surface, $f:S\rightarrow S$ an area preserving orientation preserving $C^{1}$ diffeomorphism of $S$. Let $p$ be a saddle fixed point of $f$ and assume that the branches of $p$ are invariant.
  1. (1)
    Let $L$ be a branch of $p$ and suppose that all fixed points of $f$ contained in $cl_{S}L$ are non degenerate. Then either $L$ is a connection or $L$ accumulates on both sectors adjacent to itself.
  2. (2)
    Let $L_{1}$ and $L_{2}$ be adjacent branches of $p$ that are not connections and suppose that all fixed points of $f$ contained in $cl_{S}(L_{1}\cup L_{2})$ are non degenerate. Then $cl_{S}L_{1}=cl_{S}L_{2}$.
  3. (3)
    Assume that the four branches of $p$ are not connections and that all fixed points of $f$ contained in $cl_{S}(W_{p}^{u}\cup W_{p}^{s})$ are non degenerate. Then the four branches of $p$ have the same closure in $S$.
Proof.
We start with the proof of item $(1)$. Let $(x,y)$ be continuous coordinates in a neighborhood $V$ of $p$, with $p$ at the origin and where $f(x,y)=(\lambda x,\lambda^{-1}y)$, where $\lambda>1$. We may assume that $L$ is the branch that contains the local branch $\{(x,y)\in V\hskip 1.42262pt|\hskip 1.42262ptx>0,y=0\}$.
First let’s prove that if $L$ does not accumulate on one of its adjacent sectors then it is a connection. Suppose that $L$ does not accumulate on, say, $\Sigma_{1}=\{(x,y)\in V\hskip 1.42262pt|\hskip 1.42262ptx>0,y>0\}$. So, making $V$ smaller if necessary, we can assume that $cl_{S}L\cap\Sigma_{1}=\varnothing$. Let $U$ be the component of $S\setminus cl_{S}L$ that contains $\Sigma_{1}$. Obviously $U$ is invariant.
Since $cl_{S}L$ is compact and connected, by Theorem 2.14 we have that $fr_{S}U$ has finitely many connected components and $U$ has a finite number ideal boundary points.
If $\beta:(0,1)\rightarrow U$ is defined by $\beta(t)=(t,t)$, then as $t\rightarrow 0^{+}$ we have that $\beta(t)\rightarrow p$ in $S$, $\beta(t)\rightarrow e$ for some $e\in E(U)$ and $\beta(t)\rightarrow b$, for some ideal boundary point of $U$. Since $(f\circ\beta)(t)\rightarrow e$ as $t\rightarrow 0^{+}$, we have that $e$ is a fixed point of $\hat{f}:C(b)\rightarrow C(b)$.
By item $(4)$ of Theorem 3.1 we have that $L$ is a connection contained in $Z(b)$. Therefore we have just proven that if $L$ does not accumulate on one of its adjacent sectors then it is a connection, or equivalently, if $L$ is not a connection then it accumulates on both adjacent sectors.
This proves item $(1)$.
In order to prove $(2)$ we will assume that $L_{1}$ contains $\{(x,y)\in V\hskip 1.42262pt|\hskip 1.42262ptx>0,y=0\}$, $L_{2}$ contains $\{(x,y)\in V\hskip 1.42262pt|\hskip 1.42262ptx=0,y>0\}$ and $\Sigma_{1}$ is as defined as before.
Since $L_{1}$ accumulates on $\Sigma_{1}$, from the dynamics inside $V$ we see that $L_{1}$ accumulates on a point of a fundamental domain of $L_{2}$. From Proposition 3.8 we know that $L_{2}\cap cl_{S}{L_{1}}$ implies that $L_{2}\subset cl_{S}{L_{1}}$.
Similarly $L_{1}\subset cl_{S}{L_{2}}$, which proves item $(2)$.
Item (3) follows from $(2)$. ∎
Item $(1)$ provides a strong dichotomy between recurring and non-recurring behavior of a branch $L$. Dynamically, if a branch is a connection then its limit set consists of one point. Otherwise, $L$ accumulates on its saddle, on its adjacent sectors and branches, and accumulates all over itself.
4.2. Accumulation implies intersection in genus zero
A homoclinic point of a saddle $p$ is a point in $W_{p}^{u}\cap W_{p}^{s}$ different from $p$.
4.2 Proposition.
Let $S$ be the two dimensional sphere $S^{2}$ and $f:S\rightarrow S$ an area preserving orientation preserving $C^{1}$ diffeomorphism of $S$. Let $p$ be a saddle fixed point of $f$ and assume that the branches of $p$ are invariant.
  1. (1)
    Assume that $L_{1}$ and $L_{2}$ are adjacent branches of $p$ that are not connections and that all fixed points of $f$ contained in $cl_{S}(L_{1}\cup L_{2})$ are non degenerate.
    Then $L_{1}\cap L_{2}\neq\varnothing$.
  2. (2)
    Assume that all branches of $p$ are not connections and that all fixed points of $f$ contained in $cl_{S}(W_{p}^{u}\cup W_{p}^{s})$ are non degenerate.
    Then all pairs of adjacent branches of $p$ intersect.
Proof.
We are going to assume that $L_{1}\cap L_{2}=\varnothing$ and get a contradiction.
Let $(x,y)$ be continuous coordinates in a neighborhood $V$ of $p$, with $p$ at the origin and where $f(x,y)=(\lambda x,\lambda^{-1}y)$, $\lambda>1$.
Let $L_{1}$ be the unstable branch that contains the local branch $\{(x,y)\in V\mid y=0,x>0\}$ and $L_{2}$ be the stable branch that contains the local branch $\{(x,y)\in V\mid x=0,y>0\}$. From item $(1)$ of Proposition 4.1, we have that $L_{1}$ and $L_{2}$ accumulate on the sector $\{(x,y)\in V\thinspace|\thinspace x>0,y>0\}$. We choose
$\Sigma_{0}=\{(x,y)\in V\mid 0<xy<\lambda^{-2},0<x<1,0<y<1\}$.
to represent it. Let
$\Sigma=\{(x,y)\in V\mid 0<xy\leq\lambda^{-2},0<x\leq 1,0<y\leq 1\}$,
$En(\Sigma)=\{(x,y)\in\Sigma\mid\lambda^{-1}<y\leq 1\}$ and
$Ex(\Sigma)=\{(x,y)\in\Sigma\mid\lambda^{-1}<x\leq 1\}$.
Note that $En(\Sigma)$ and $Ex(\Sigma)$ are disjoint.
From the dynamics of $f$ in $\Sigma$, we see that every orbit of $f$ that visits $\Sigma$ must enter $\Sigma$ through $En(\Sigma)$, meaning that if $x\notin\Sigma$ and $f^{n}(x)\in\Sigma$ for some positive $n$, then for the smallest $n$ with this property, $f^{n}(x)\in En(\Sigma)$. In the same way, every orbit of $f^{-1}$ that visits $\Sigma$ must enter $\Sigma$ through $Ex(\Sigma)$.
If we start at $p$ and move along the unstable branch $L_{1}$, let $q_{1}$ be the first point of $L_{1}$ to intersect $\Sigma$. Since $L_{1}$ is unstable we have that $q_{1}\in En(\Sigma)$. Join $q_{1}$ and $p$ by a line segment $\gamma_{1}\subset\Sigma_{0}$ and consider $\Gamma_{1}=L_{1}[p,q_{1}]\cup\gamma_{1}$ where $L_{1}[p,q_{1}]$ is the segment of $L_{1}$ between $p$ and $q_{1}$. Since $S$ has genus zero, $\Gamma_{1}$ is a simple closed curve that separates $S$ into two open disks whose common frontier in $S$ is $\Gamma_{1}$.
In the same way, if we start at $p$ and move along the stable branch $L_{2}$, let $q_{2}$ be the first point of $L_{2}$ to intersect $\Sigma$. Since $L_{2}$ is stable we have that $q_{2}\in Ex(\Sigma)$. Join $q_{2}$ and $p$ by a line segment $\gamma_{2}\subset\Sigma_{0}$ and consider $\Gamma_{2}=L_{2}[p,q_{2}]\cup\gamma_{2}$, where $L_{2}[p,q_{2}]$ is the segment of $L_{2}$ between $p$ and $q_{2}$. $\Gamma_{2}$ is also a simple closed curve in $S$.
Note that the local branch of $L_{2}$ and $\gamma_{2}$ are in different components of $S-\Gamma_{1}$. From this we conclude that $\Gamma_{1}$ and $\Gamma_{2}$ must intersect again at a point $q$ different from $p$. Since $\Gamma_{1}\cap\gamma_{2}=\varnothing$ and $\Gamma_{2}\cap\gamma_{1}=\varnothing$ we have that $q\in L_{1}[p,q_{1}]\cap L_{2}[p,q_{2}]$, a contradiction. This proves $(1)$.
Item $(2)$ follows from $(1)$. ∎
4.3. Accumulation implies intersection in genus one
Let us now consider the case where $S$ is the two dimensional torus $T^{2}$.
4.3 Proposition.
Let $S$ be a connected compact orientable surface of genus $g$ and $f:S\rightarrow S$ an orientation preserving area preserving $C^{1}$ diffeomorphism of $S$. Let $p$ be a saddle fixed point of $f$ and assume that the branches of $p$ are invariant.
Assume that the four branches of $p$ are not connections and that all fixed points of $f$ contained in $cl_{S}(W_{p}^{u}\cup W_{p}^{s})$ are non degenerate.
Then we know the following:
  1. (1)
    If $g=1$ then $p$ has homoclinic points.
  2. (2)
    For any $g\geq 1$, if $p$ has homoclinic points then all pairs of adjacent branches of $p$ intersect.
Before the proof, a couple of remarks.
The conclusions about the existence of homoclinic points in Propositions 4.2 and 4.3 are false if the genus of the surface is greater than one. Examples could be the time one map of area preserving flows on surfaces with finitely many singularities and every other orbit dense.
Item $(1)$ of Proposition 4.2 is not true if the genus of $S$ is one. In the torus two branches could close a connection and the other two spin around the torus like a line of irrational slope without intersecting.
Proof.
We think of $S$ as $\mathbb{R}^{2}/\mathbb{Z}^{2}$.
Let $E=\mathbb{R}^{2}/(3\mathbb{Z}\times\mathbb{Z})$ and $\pi:E\rightarrow S$ the canonical 3-fold covering of the torus, whose group of deck transformations is isomorphic to $\mathbb{Z}_{3}=\mathbb{Z}/3\mathbb{Z}$ and generated by the rotation $T(x,y)=(x+1,y)$.
Let $F:E\rightarrow E$ be the lifting of $f$ that fixes the points of $\pi^{-1}(p)=\{p_{0},p_{1},p_{2}\}$. If $L$ is a branch of $p$ then $\pi^{-1}(L)=L_{0}\cup L_{1}\cup L_{2}$, where every $L_{i}$ is a branch of $p_{i}$ and $L_{i+1}=TL_{i}$, for $i\in\mathbb{Z}_{3}$.
In order to prove that $p$ has homoclinic points it is enough to show that for some pair $(i,j)$ an unstable branch of $p_{i}$ intersects a stable branch of $p_{j}$.
By Proposition 4.1, we know that the branches of $p$ have the same closure in $S$. Therefore the same happens with the points $p_{i}$. Let $K_{i}$ be the closure in $E$ of the branches of each $p_{i}$. We have that $T(K_{i})=K_{i+1}$ for $i\in\mathbb{Z}_{3}$.
Let =$\mathcal{I}=\{(i,j)\in\mathbb{Z}_{3}\times\mathbb{Z}_{3}\mid i\neq j\}$. We have two possibilities:
  1. (1)
    There exists $(i,j)\in\mathcal{I}$ such that $p_{i}\in K_{j}$.
  2. (2)
    For every $(i,j)\in\mathcal{I}$ there exists a neighborhood $V_{ij}$ of $p_{i}$ such that $V_{ij}\cap K_{j}=\varnothing$.
4.4 Lemma.
If there exists $(i,j)\in\mathcal{I}$ such that $p_{i}\in K_{j}$ then $K_{0}=K_{1}=K_{2}$.
Proof.
We assume that $(i,j)=(2,1)$, that is, $p_{2}\in K_{1}$. The argument with other elements of $\mathcal{I}$ is the same.
By Proposition 3.12 we have that $W_{p_{2}}^{u}\cup W_{p_{2}}^{s}\subset K_{1}$. Therefore $K_{2}\subset K_{1}$. It follows that $K_{0}=T(K_{2})\subset T(K_{1})=K_{2}$ and $K_{1}=T(K_{0})\subset T(K_{2})=K_{0}$. Hence $K_{0}=K_{1}=K_{2}$. ∎
4.5 Lemma.
If for every $(i,j)\in\mathcal{I}$ there exists a neighborhood $V_{ij}$ of $p_{i}$ such that $V_{ij}\cap K_{j}=\varnothing$, then for every $i$ there exists a neighborhood $V_{i}$ of $p_{i}$ such that $V_{i}\cap K_{j}=\varnothing$ for every $j\neq i$.
Proof.
$V_{1}=V_{12}\cap V_{13}$ is a neighborhood of $p_{1}$ disjoint from $K_{2}$ and $K_{3}$. In the same way we construct $V_{2}$ and $V_{3}$. ∎
From Lemmas 4.4 and 4.5, we conclude that we have the following possibilities:
  1. (1)
    $K_{0}=K_{1}=K_{2}$.
  2. (2)
    For every $i$ there exists a neighborhood $V_{i}$ of $p_{i}$ such that $V_{i}\cap K_{j}=\varnothing$ for every $j\neq i$.
Let us consider the case $K_{0}=K_{1}=K_{2}$. The argument will be by contraction. So let us assume that $p$ has no homoclinic points.
Using items $(1)$ and $(3)$ of Proposition 4.1, it is easy to conclude that for every $i$ and $j$, each branch of $p_{i}$ accumulates on the four sectors of $p_{j}$.
Let $(x,y)$ be continuous coordinates in a neighborhood $V$ of $p_{1}$, with $p_{1}$ at the origin and where $f(x,y)=(\lambda x,\lambda^{-1}y)$, $\lambda>1$.
We are going to consider the sector represented by
$\Sigma_{0}=\{(x,y)\in V\mid 0<xy<\lambda^{-2},0<x<1,0<y<1\},$
and the sets $\Sigma$, $En(\Sigma)$ and $Ex(\Sigma)$ as defined in the proof of Proposition 4.2.
We are going to denote the unstable branches of $p_{i}$ by $W_{+}^{u}(p_{i})$ and $W_{-}^{u}(p_{i})$ and stable ones by $W_{+}^{s}(p_{i})$ and $W_{-}^{s}(p_{i})$. All branches of all points $p_{i}$ accumulate on $\Sigma_{0}$.
Let $q_{+}^{u}(p_{i})\in W_{+}^{u}(p_{i})$ and $q_{-}^{u}(p_{i})\in W_{-}^{u}(p_{i})$ be the first point of each unstable branch to intersect $\Sigma$ as we move along the corresponding branch starting from $p_{i}$. Both $q_{+}^{u}(p_{i})$ and $q_{-}^{u}(p_{i})$ belong to $En(\Sigma)$. Join $q_{+}^{u}(p_{i})$ and $q_{-}^{u}(p_{i})$ by a small arc $\gamma_{i}^{u}$ contained in $En(\Sigma)$. Let $W^{u}[q_{-}^{u}(p_{i}),q_{+}^{u}(p_{i})]$ be the segment inside $W^{u}(p_{i})$ from $q_{-}^{u}(p_{i})$ to $q_{+}^{u}(p_{i})$. Then we have that $\Gamma_{i}^{u}=W^{u}[q_{-}^{u}(p_{i}),q_{+}^{u}(p_{i})]\cup\gamma_{i}^{u}$ is a simple closed curve that contains $p_{i}$.
In the same way let $q_{+}^{s}(p_{i})\in W_{+}^{s}(p_{i})$ and $q_{-}^{s}(p_{i})\in W_{-}^{s}(p_{i})$ be the first point of each stable branch to intersect $\Sigma$ as we move along the corresponding branch starting from $p_{i}$. Both $q_{+}^{s}(p_{i})$ and $q_{-}^{s}(p_{i})$ belong to $Ex(\Sigma)$. Join $q_{+}^{s}(p_{i})$ and $q_{-}^{s}(p_{i})$ by a small arc $\gamma_{i}^{s}$ contained in $En(\Sigma)$. Let $W^{s}[q_{-}^{s}(p_{i}),q_{+}^{s}(p_{i})]$ be the segment inside $W^{s}(p_{i})$ from $q_{-}^{s}(p_{i})$ to $q_{+}^{s}(p_{i})$. Then we have that $\Gamma_{i}^{s}=W^{s}[q_{-}^{s}(p_{i}),q_{+}^{s}(p_{i})]\cup\gamma_{i}^{s}$ is a simple closed curve that contains $p_{i}$.
We have that $\gamma_{i}^{u}\cap\gamma_{j}^{s}=\varnothing$ for all pairs $(i,j)$. Therefore $\Gamma_{i}^{u}\cap\Gamma_{j}^{s}=\varnothing$ if $i\neq j$ and $\Gamma_{i}^{u}\cap\Gamma_{i}^{s}=\{p_{i}\}$.
If $\alpha$ and $\beta$ are closed curves in $T^{2}$ let #$\#(\alpha,\beta)$ be the oriented intersection number of $\alpha$ and $\beta$. ##\## depends only on the homology classes of $\alpha$ and $\beta$ and is a non degenerate skew symmetric bilinear form over the integers defined on $H_{1}(T^{2},\mathbb{Z})\sim\mathbb{Z}^{2}$.
We have $E$ is homeomorphic to $T^{2}$ and if we look at $\Gamma_{i}^{u}$ and $\Gamma_{i}^{s}$ as elements of $H_{1}(T^{2},\mathbb{Z})$ then |#$|\#(\Gamma_{i}^{u},\Gamma_{j}^{s})|=\left\{\begin{array}[]{c}1\text{ if }i=j\text{}\\ 0\text{ if }i\neq j\text{}\text{}\text{ }\end{array}\right.$.
The rank of $H_{1}(T^{2},\mathbb{Z})$ is $2$ and therefore $\{\Gamma_{1}^{s},\Gamma_{2}^{s},\Gamma_{3}^{s}\}$ is linearly dependent. So one of them is a linear combination of the other two, say $\Gamma_{1}^{s}=a\Gamma_{2}^{s}+b\Gamma_{3}^{s}$. It follows that $01=|\#(\Gamma_{1}^{u},\Gamma_{1}^{s})|=|a||\#(\Gamma_{1}^{u},\Gamma_{2}^{s})|+|b||\#(\Gamma_{1}^{u},\Gamma_{3}^{s})|=$ 0, a contradiction.
This proves that $p$ has homoclinic points if $K_{0}=K_{1}=K_{2}$.
Now let us consider the case where for every $i$ there exists a neighborhood $V_{i}$ of $p_{i}$ such that $V_{i}\cap K_{j}=\varnothing$ for every $j\neq i$. We may assume that the $V_{0}$, $V_{1}$ and $V_{2}$ are pairwise disjoint.
The procedure is almost the same. We start by assuming by contradiction that $p$ does not have homoclinic points. The difference in this case is that the branches of each $p_{i}$ only accumulate on the sectors of $p_{i}$, not on the sectors of $p_{j}$ if $j\neq i$.
For each $p_{i}$ we take a sector $\Sigma_{0}(i)\subset V_{i}$ and use it to construct a pair of simple closed curves ($\Gamma_{i}^{u}$ , $\Gamma_{i}^{s}$) that satisfy $\Gamma_{i}^{u}\cap\Gamma_{j}^{s}=\varnothing$ if $i\neq j$ and $\Gamma_{i}^{u}\cap\Gamma_{i}^{s}=\{p_{i}\}$. The contradiction is obtained in the same way.
This completes the proof of item (1) that $p$ has a homoclinic point.
Now, we would like to prove that for any $g\geq 1$, if $p$ has homoclinic points then all pairs of adjacent branches of $p$ intersect.
We need the following result:
4.6 Lemma.
Let $M$ be a compact orientable surface of genus $g$ and $\alpha_{0},...,\alpha_{2g}$ be a collection of $12g+$ 1 closed curves of $M$ such that for every $i=0,...,2g$ we have that $\alpha_{i}\setminus\underset{j\neq i}{\bigcup}\alpha_{j}\neq\varnothing$. Then $M\setminus\underset{0\leq i\leq 2g}{\bigcup}\alpha_{i}$ is disconnected.
Proof.
Assume by contradiction that $M\setminus\underset{0\leq i\leq 2g}{\bigcup}\alpha_{i}$ is connected. For $0\leq i\leq 2g$, let $x_{i}\in\alpha_{i}\setminus\underset{j\neq i}{\bigcup}\alpha_{j}$ and $D_{i}$ a small disk containing $x_{i}$ and disjoint from $\underset{j\neq i}{\bigcup}\alpha_{j}$.
Let $y_{i}$ and $z_{i}$ be points of $D_{i}$ on different sides of $\alpha_{i}$. Let $\beta_{1i}$ be a small curve connecting $y_{i}$ and $z_{i}$ inside $D_{i}$. Since $M\setminus\underset{0\leq i\leq 2g}{\bigcup}\alpha_{i}$ is connected, there exists a simple curve $\beta_{2i}$ connecting $z_{i}$ to $y_{i}$ inside $M\setminus\underset{0\leq i\leq 2g}{\bigcup}\alpha_{i}$. If $\beta_{i}=\beta_{1i}*\beta_{2i}$, then $\beta_{i}$ is a simple closed curve with the following properties:
  1. (1)
    The oriented intersection number #$\#(\alpha_{i},\beta_{i})=1$, if we choose suitable orientations for $M$, $\alpha_{i}$ and $\beta_{i}$.
  2. (2)
    #$\#(\alpha_{j},\beta_{i})=0$ for $j\neq i$, since $\beta_{i}\cap\big{(}\underset{j\neq i}{\bigcup}\alpha_{j}\big{)}=\varnothing$
Let $h_{i}:H_{1}(M,\mathbb{R})\rightarrow\mathbb{R}$ be the homomorphism $h_{i}(\gamma)=\#(\gamma,\beta_{i})$. It follows from $(1)$ and $(2)$ that $h_{0},...,h_{2g}$ are linearly independent elements of $\big{(}H_{1}(M,\mathbb{R})\big{)}^{*}$, contradicting the fact that the dimension of $\big{(}H_{1}(M,\mathbb{R})\big{)}^{*}$ is $2g$.
Returning to the proof of item (2) of Proposition 4.3, we know that two branches of $p$, say $W_{+}^{u}$ and $W_{+}^{s}$, intersect at a homoclinic point $q$.
Let $\alpha_{0}=W_{+}^{u}[q,f(q)]\cup W_{+}^{s}[q,f(q)]$. $\alpha_{0}$ is a closed curve.
For $0\leq i\leq 2g$, consider the closed curves
$\alpha_{i}=f^{i}({\alpha_{0})}=W_{+}^{u}[f^{i}(q),f^{i+1}(q)]\cup W_{+}^{s}[f^{i}(q),f^{i+1}(q)]$ and $\Omega=\bigcup_{0\leq i\leq 2g}\alpha_{i}$.
We would like to show that, since $p$ has no connections, the curves $\alpha_{0}$,…,$\alpha_{2g}$ satisfy the hypothesis of Lemma 4.6.
There exists a point $\bar{q}\in W_{+}^{u}$ such that $\bar{q}\notin W_{+}^{s}$. We have that $o(\bar{q})\subset W_{+}^{u}$ and $o(\bar{q})\cap W_{+}^{s}=\varnothing$. We may assume that $\bar{q}\in W_{+}^{u}(q,f(q))$. For every $k\in\mathbb{Z}$ we have that $f^{k}\bar{q}\in W_{+}^{u}(f^{i}(q),f^{i+1}(q)$ if and only if $k=i$. Since $f^{k}\bar{q}\notin W_{+}^{s}$, for every $k\in{0,...,2g}$ we have that $f^{k}\bar{q}\in\alpha_{k}\setminus\underset{0\leq j\leq 2g,j\neq k}{\bigcup}\alpha_{j}$, and $\alpha_{0}$,…,$\alpha_{2g}$ satisfy the hypothesis of Lemma 4.6.
It follows that $S\setminus\Omega$ is disconnected. Observe that $\Omega=\Omega_{u}\cup\Omega_{s}$, where
$\Omega_{u}=W_{+}^{u}[q,f^{2g+1}(q)]$ and $\Omega_{s}=W_{+}^{s}[q,f^{2g+1}(q)].$
Since $\Omega_{u}$ and $\Omega_{s}$ are simple arcs homeomorphic to a closed interval, we have that the frontier in $S$ of every component of $S\setminus\Omega$ has points in $\Omega_{u}\subset W_{+}^{u}$ and in $\Omega_{s}\subset W_{+}^{s}$.
Let $A$ be the component of $S\setminus\Omega$ that contains $p$ and its local branches, and $B$ any other component. By recurrence, there exist infinitely many values of $n$ such that $f^{n}B\cap B\neq\varnothing$.
Since $fr_{S}B\cap W_{+}^{s}\neq\varnothing$, we have that $f^{n}B\cap A\neq\varnothing$ for every $n$ large enough. Hence $f^{n}B\cap fr_{S}A\neq\varnothing$, implying that $B$ contains points of $W_{+}^{u}$ or $W_{+}^{s}$.
We have that the four branches have the same closure in $S$. This forces $W_{-}^{u}$ and $W_{-}^{s}$ to leave $A$ and enter $B$, by intersecting $fr_{S}A$ which is contained in $W_{+}^{u}\cup W_{+}^{s}$. Therefore $W_{-}^{u}$ must intersect $W_{+}^{s}$ and $W_{-}^{s}$ must intersect $W_{+}^{u}$.
It remains to show that $W_{-}^{u}\cap W_{-}^{s}\neq\varnothing$. But this can be achieved by repeating the argument with a homoclinic intersection of $W_{+}^{u}$ and $W_{-}^{s}$ or of $W_{-}^{u}$ and $W_{+}^{s}$. ∎
5. Real analytic diffeomorphisms and the standard map
The standard map is a one parameter family of area preserving diffeomorphism of the two dimensional torus $T^{2}=\mathbb{R}^{2}/\mathbb{Z}^{2}$ given by
$f_{\mu}(x,y)=(x+y+\frac{\mu}{2\pi}\sin(2\pi x),y+\frac{\mu}{2\pi}\sin(2\pi x))$, $\mu\in\mathbb{R}$.
$f_{0}$ is just a linear twist. $\varphi(x,y)=(\frac{1}{2}-x,-y)$ is a conjugacy between $f_{\mu}$ and $f_{-\mu}$, so we only consider parameters $\mu>0$.
For $\mu\neq 0$ there are two fixed points, $p=(0,0)$ and $q=(\frac{1}{2},0)$. $p$ is always a saddle with positive eigenvalues and it is called the principal fixed point of $f_{\mu}$. $q$ is elliptic if $40<\mu<$ 4 and a hyperbolic fixed point with negative eigenvalues if $\mu>4$.
In this section we are going to prove the following.
5.1 Theorem.
If $\mu\neq 4$ then the standard map satisfies the following:
  1. (1)
    The four branches of $p=(0,0)$ have topologically transverse homoclinic points.
  2. (2)
    $f_{\mu}$ has positive topological entropy.
  3. (3)
    There exist periodic saddles with transverse homoclinic points.
5.1. The principal point $p$ always has homoclinic points
For $\mu\neq 4$, if $L$ is a branch of $p$ then $L$ is invariant and all fixed points of $f$ contained in $cl_{S}L$ are non degenerate.
Here we make use of the weaker hypothesis that non degeneracy refers only to fixed points of $f$ in $cl_{S}L$.
We have that $f(-x,-y)=-f(x,y)$ implying that the invariant manifolds of $p$ are symmetric with respect to $(0,0)$. Therefore if $L$ is a connection then so is $-L$.
There can not be a connection between $p$ and $q$. This is obvious when $q$ is elliptic. When $q$ is a hyperbolic this can not happen because the branches of $p$ are invariant and those of $q$ have period two.
Therefore, if one of the branches of $p$ is a connection, then that connection is equal to two branches of $p$. By symmetry, the invariant manifolds of $p$ are formed of two connections of homoclinic points.
If no branch of $p$ is a connection then by Proposition 4.3 the four branches of $p$ still have homoclinic points. Therefore we have the following:
5.2 Proposition.
For $\mu\neq 4$, either the four branches of $p$ form two connections symmetric with respect $(0,0)$, or no branch of $p$ is a connection and each has homoclinic points.
Next, we are going to show that the first alternative can not happen.
5.2. No branch of $p$ is a connection
Now we present Ushiki’s theorem.
Let $F_{\mu}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$, $F_{\mu}(x,y)=(x+y+\frac{\mu}{2\pi}\sin(2\pi x),y+\frac{\mu}{2\pi}\sin(2\pi x))$, be the lifting of $f_{\mu}$ to the universal cover. Since a connection $L$ of $p=(0,0)$ in $\mathbb{T}^{2}$ connects $p$ to itself, a lifting of $L$ would be a connection $\widehat{L}$ connecting two points of $\mathbb{Z}^{2}$.
5.3 Proposition.
(Ushiki’s theorem) If a real analytic diffeomorphism $F:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ extends into an automorphism $H:\mathbb{C}^{2}\rightarrow\mathbb{C}^{2}$ of the complex space of dimension two, then $F$ has no connections.
This is Theorem $1$ of [16].
It follows from basic classification results of complex $1$-manifolds.
Let $H_{\mu}:\mathbb{C}^{2}\rightarrow\mathbb{C}^{2}$,
$H_{\mu}(z_{1},z_{2})=(z_{1}+z_{2}+\frac{\mu}{2\pi}\sin(2\pi z_{1}),z_{2}+\frac{\mu}{2\pi}\sin(2\pi z_{1}))$
be an extension of $F_{\mu}$ to $\mathbb{C}^{2}$. If we write $(w_{1},w_{2})=H_{\mu}(z_{1},z_{2})$, then a simple calculation shows that
$\begin{cases}{z_{1}}={w_{1}}-{w_{2}}\\ z_{2}=w_{2}-\frac{\mu}{2\pi}{\sin\big{(}2\pi(w_{1}-w_{2})\big{)}}.\end{cases}$
It follows that $H_{\mu}$ is an automorphism of $\mathbb{C}^{2}$, and from Proposition 5.3 we have that $F_{\mu}$ has no connections.
This proves the following.
5.4 Proposition.
For $\mu\neq 4$, $f_{\mu}$ can not have connections.
5.3. Every branch $p$ has a topologically transverse homoclinic point
In this subsection we are going to prove the following.
5.5 Proposition.
For $\mu\neq 4$, the four branches of the principal point $p$ have topologically transverse homoclinic points.
More precisely, for every pair of adjacent branches $L^{u}$, $L^{s}$ of $p$ there exist $q\in L^{u}\cap L^{s}$ and a small open disk $D$ containing $q$ such that, if $\gamma_{u}$ and $\gamma_{s}$ are the components of $L^{u}\cap D$ and $L^{s}\cap D$ that contain $q$, respectively, then they satisfy the following:
  1. (1)
    $\gamma_{u}\cap\gamma_{s}=\{q\}$.
  2. (2)
    $D\setminus\gamma_{u}$ and $D\setminus\gamma_{s}$ have two components.
  3. (3)
    The components of $\gamma_{u}\setminus\{q\}$ are contained in different components of $D\setminus\gamma_{s}$.
  4. (4)
    The components of $\gamma_{s}\setminus\{q\}$ are contained in different components of $D\setminus\gamma_{u}$.
We know from item $(2)$ if Proposition 4.3 that all pairs of adjacent branches of $p$ intersect.
We are going to work the case where we start from $q\in W_{+}^{u}\cap W_{+}^{s}$, and then show that the pairs ($W_{-}^{s}$, $W_{+}^{u}$) and ($W_{-}^{u}$, $W_{+}^{s}$) have a topologically transverse intersection. Then repeating the argument starting with $q\in W_{-}^{u}\cap W_{+}^{s}$ we show that the other two pairs ($W_{-}^{s}$, $W_{-}^{u}$) and ($W_{+}^{u}$, $W_{+}^{s}$) have a topologically transverse intersection.
We are only going to prove that ($W_{-}^{s}$, $W_{+}^{u}$) have a topologically transverse intersection, the argument for ($W_{-}^{u}$, $W_{+}^{s}$) is analogous.
Let $\Omega_{u}=W_{+}^{u}[q,f^{3}q]$, $\Omega_{s}=W_{+}^{s}[q,f^{3}q]$ and $\Omega=\Omega_{u}\cup\Omega_{s}$.
Since $p$ has no connections, we have that $T^{2}\setminus\Omega$ is disconnected by Lemma 4.6.
The frontier in $S$ of every component of $S\setminus\Omega$ is contained in $\Omega$ and intersects both $\Omega_{u}\subset W_{+}^{u}$ and $\Omega_{s}\subset W_{+}^{s}$.
Let $A$ be the component of $T^{2}\setminus\Omega$ that contains $p$. We saw that the local branch of $W_{-}^{s}$ must leave $A$, intersect $\Omega$ (at at least one point) and then enter a different component of $T^{2}-\Omega$. Let $C$ be the union of all components of $T^{2}\setminus\Omega$ different from $A$.
We denote the eigenvalues of $df_{p}$ by $\lambda>1$ and $\lambda^{-1}$. There exist real analytic immersions
[…]
such that
  1. (1)
    $f\big{(}\alpha(t)\big{)}=\alpha(\lambda t)\hskip 4.2679pt\forall t\in\mathbb{R}$, $\alpha(0)=p$ and $df_{p}\big{(}\alpha^{\prime}(0)\big{)}=\lambda\thinspace\alpha^{\prime}(0)$.
  2. (2)
    $f\big{(}\beta(s)\big{)}=\beta(\lambda^{-1}s)\hskip 4.2679pt\forall s\in\mathbb{R}$, $\beta(0)=p$ and $df_{p}\big{(}\beta^{\prime}(0)\big{)}=\lambda^{-1}\thinspace\beta^{\prime}(0)$.
We may assume that $\beta(0,+\infty)=W_{-}^{s}$. Note that $\beta(0,+\infty)$ can not intersect $\Omega_{s}\subset W_{+}^{s}$. Since all pairs of adjacent branches of $p$ intersect, we have that $\beta(0,+\infty)$ intersects only $\Omega_{u}$.
For every positive and sufficiently small $s$, we have that $\beta\big{(}[0,s]\big{)}\subset A$. We also know that $\beta(s)\in C$ for some values of $s>0$. Therefore the set
$E=\{s\geq 0\mid\beta\big{(}[0,s]\big{)}\subset A\cup\Omega_{u}\}$
is not empty and bounded above. Let $\bar{s}:=supE$. It easy to see that $E$ is the interval $[0,\bar{s}]$ and that $\beta(\bar{s})\in\Omega_{u}$.
We will need the following.
5.6 Lemma.
There exists $\varepsilon>0$ such that $\beta(\overline{s}-\varepsilon,\overline{s})\subset A$ and $\beta(\overline{s},\overline{s}+\varepsilon)\subset C$.
Proof.
We are going to work with the lift to $\mathbb{R}^{2}$ of various objects we are dealing with. We will use the same notation for an object downstairs and for its lift to the universal cover.
Firstly, we will make a real analytical change of coordinates at $p=(0,0)$, so that the local invariant manifolds also become the coordinate axes. Consider the mapping $\Phi:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$, given by $\Phi(t,s)=\alpha(t)+\beta(s)$. We have that
$\Phi^{\prime}(0,0)=\begin{bmatrix}\alpha_{1}^{\prime}(0)&\beta_{1}^{\prime}(0)\\ \alpha_{2}^{\prime}(0)&\beta_{2}^{\prime}(0)\\ \end{bmatrix},$
and therefore $det\hskip 1.42262pt\Phi^{\prime}(0,0)\neq 0$, since $\alpha^{\prime}(0)$ and $\beta^{\prime}(0)$ are eigenvectors of $df_{p}$ corresponding to different eigenvalues.
There exist neighborhoods $\mathcal{U}$ and $\mathcal{V}$ of $(0,0)$ such that the restriction $\Phi:\mathcal{U}\rightarrow\mathcal{V}$ is a real-analytic diffeomorphism. Its inverse is also real-analytic diffeomorphism, and will be denoted by $\Psi:\mathcal{V}\rightarrow\mathcal{U}$. We will also assume that $\mathcal{U}$ is the ball $B_{\delta}$ with center at $p$ and radius $\delta$, where $\delta$ is small.
The local invariant manifolds defined by $\mathcal{V}$, $W_{loc}^{u}$ and $W_{loc}^{s}$, are the components of $W^{u}\cap\mathcal{V}$ and $W^{s}\cap\mathcal{V}$, respectively, that contain $p=(0,0)$.
We have that $\Psi\big{(}\alpha(t)+\beta(s)\big{)}=(t,s),\thinspace\forall\thinspace(t,s)\in B_{\delta}$. Therefore, for $t,s\in(-\delta,\delta)$ we have that
$\Psi$($\alpha(t)\big{)}=(t,0)$    and    $\Psi$($\beta(s)\big{)}=(0,s)$,
meaning that $\alpha$ and $\beta$ could be thought as coordinate curves in the original coordinates. If $\Psi_{i}$, $i=1,2$, are the coordinate functions of $\Psi$ then
$W_{loc}^{u}=\{(x,y)\in\mathcal{V}\thinspace|\thinspace\Psi_{2}(x,y)=0\}$ and $W_{loc}^{s}=\{(x,y)\in\mathcal{V}\thinspace|\thinspace\Psi_{1}(x,y)=0\}$.
We may assume that $f^{i}q\in W_{+loc}^{u}$ for $0\leq i\leq 3$ and therefore $\Omega_{u}\subset W_{+loc}^{u}$. It follows that $\Psi(\Omega_{u})\subset\{(t,s)\in B_{\delta}\mid s=0\}$.
We have that $(\Psi_{2}\circ\beta)(\bar{s})=0$. Therefore there exists $\delta>0$ such that $\beta(\bar{s}-\delta,\bar{s}+\delta)\subset\mathcal{V}$. It follows that the composition $\Psi_{2}\circ\beta:(\bar{s}-\delta,\bar{s}+\delta)\rightarrow\mathbb{R}$ is a well defined real analytic function.
For $s\in(\bar{s}-\delta,\bar{s}+\delta)$ we have that $(\Psi_{2}\circ\beta)(s)=0$ if and only if $\beta(s)\in\Omega_{u}$.
Since the zeroes of a non constant analytic function are isolated, we know that there exists $\varepsilon\in(0,\delta)$ such that $\Psi_{2}\circ\beta:(\bar{s}-\varepsilon,\bar{s}+\varepsilon)\rightarrow\mathbb{R}$ satisfies one of the following conditions.
  1. (1)
    $(\Psi_{2}\circ\beta)(s)=0$ for every $s\in(\bar{s}-\varepsilon,\bar{s}+\varepsilon)$.
  2. (2)
    If $s\in(\bar{s}-\varepsilon,\bar{s}+\varepsilon)$ then $(\Psi_{2}\circ\beta)(s)=0$ if and only if $s=\bar{s}$.
Since $\bar{s}=sup{E}$, we know that for every sufficiently large $n$ there exists $s_{n}$ such that $\bar{s}<s_{n}<\bar{s}+\frac{1}{n}$ and $\beta(s_{n})\in C$. This implies that $\beta(s_{n})\notin\Omega_{u}$ and $(\Psi_{2}\circ\beta)(s_{n})\neq 0$ for every $n$. From this we conclude that condition $(1)$ can not happen.
From condition $(2)$ we have that if $s\in(\bar{s}-\varepsilon,\bar{s})$ or $s\in(\bar{s},\bar{s}+\varepsilon)$, then $\Psi_{2}(\beta(s))\neq 0$, which implies that $\beta(s)\notin\Omega_{u}$. If $s\in(\bar{s}-\varepsilon,\bar{s})$ then $\beta(s)\in A$ and if $s\in(\bar{s},\bar{s}+\varepsilon)$ then $\beta(s)\in C$.
This proves Lemma 5.6. ∎
Let $D$ be a disk that contains $q:=\beta(\bar{s})$ and is small enough so that each of the arcs $\beta(\bar{s}-\varepsilon,\bar{s}+\varepsilon)$ and $\Omega_{u}$ separate $D$ into two components. Let $\gamma_{s}$ and $\gamma_{u}$ be the components of $D\cap\beta(\bar{s}-\varepsilon,\bar{s}+\varepsilon)$ and $D\cap\Omega_{u}$, respectively, which contain $q$.
Then it is easy to verify that $D$, $\gamma_{s}$ and $\gamma_{u}$ satisfy the conditions $(1)$ to $(4)$ of Proposition 5.5.
This proves item $(1)$ of Theorem 5.1.
5.4. Positive topological entropy and transverse homoclinic points
Topologically transverse homoclinic points force the dynamics to have positive topological entropy.
Let $f$ be a $C^{1}$ diffeomorphism of a surface $S$.
5.7 Proposition.
Let $p$ be a hyperbolic periodic point of $f$ of saddle type. Assume that two branches of $p$ have a topologically transverse homoclinic point.
Then some power of $f$ has the full shift on two symbols as a topological factor, i.e., there is a subset $\Lambda$ that is invariant under $f^{n}$ for some $n>1$, and a continuous map $\pi:\Lambda\rightarrow\Sigma$, that is onto but not necessarily injective, such that $\pi\circ f^{n}=\sigma\circ\pi$, where $\sigma:\Sigma\rightarrow\Sigma$ is the full shift on two symbols.
We would like to point that the topological crossing could be of infinite order.
See Theorems $2.1$, $2.4$ and Lemma $2.8$ of [2] for a proof. The idea boils down to finding a rectangle and strips that are stretched inside the rectangle, as in the usual horseshoe.
From basic properties of topological entropy we have
$h_{top}(f)\geq\frac{1}{n}h_{top}(f^{n}|\Lambda)\geq\frac{1}{n}h_{top}(\sigma)>0$.
Therefore we have proved item $(2)$ of Theorem 5.1.
5.8 Proposition.
For $\mu\neq 4$, the standard map has positive topological entropy.
A well-known result of Katok, shows that in the case of $C^{1+\alpha}$ surface diffeomorphisms, the positiveness of the entropy is equivalent to the existence of transverse homoclinic points.
From this we have item $(3)$ of Theorem 5.1.
5.9 Proposition.
For $\mu\neq 4$, the standard map has saddles with transverse homoclinic points.
This completes the Proof of Theorem 5.1.
5.5. Positive topological entropy for real analytic diffeomorphisms
Finally, we would like to remark that the arguments presented in this section can be used to prove the following result.
5.10 Proposition.
Let $S$ be a connected orientable real analytic surface of finite genus $g$ and $f:S\rightarrow S$ be an orientation preserving real analytic diffeomorphism that preserves a finite measure $\mu$ which is positive on open sets.
Let $p$ be a saddle fixed point of $f$ that satisfies the following conditions.
  1. (1)
    $p$ has no connections.
  2. (2)
    The four branches of $p$ are relatively compact in $S$.
  3. (3)
    Every fixed point of $f$ in $cl_{S}({W_{p}^{u}\cup W_{p}^{s}})$ is non degenerate.
  4. (4)
    $p$ has homoclinic points (which happens automatically if $g=0$ or $1$).
Then $p$ has topologically transverse homoclinic points and $h_{top}(f)>0$.
The idea is to consider the ideal completion $B(S)$ of $S$, which in this case is a compact surface that contains $S$ as an open and dense subset. The ideal boundary is totally disconnected and the finite measure extends to $B(S)$. Then we do the same reasoning.
Declarations
  • Funding: not applicable.
  • Conflict of interest/Competing interests: not applicable.
  • Ethics approval: not applicable.
  • Consent to participate: not applicable.
  • Consent for publication: not applicable.
  • Availability of data and materials: not applicable.
  • Code availability: not applicable.
  • Authors’ contributions: not applicable.